### 12 Mar 2024 v3

# Group Decision-Making among Privacy-Aware Agents

Abstract.
How can individuals exchange information to learn from each other despite their privacy needs and security concerns? For example, consider individuals deliberating a contentious topic and being concerned about divulging their private experiences. Preserving individual privacy and enabling efficient social learning are both important desiderata but seem fundamentally at odds with each other and very hard to reconcile. We do so by controlling information leakage using rigorous statistical guarantees that are based on differential privacy (DP). Our agents use log-linear rules to update their beliefs after communicating with their neighbors. Adding DP randomization noise to beliefs provides communicating agents with plausible deniability with regard to their private information and their network neighborhoods. We consider two learning environments one for distributed maximum-likelihood estimation given a finite number of private signals and another for online learning from an infinite, intermittent signal stream. Noisy information aggregation in the finite case leads to interesting tradeoffs between rejecting low-quality states and making sure all high-quality states are accepted in the algorithm output. Our results flesh out the nature of the trade-offs in both cases between the quality of the group decision outcomes, learning accuracy, communication cost, and the level of privacy protections that the agents are afforded.
distributed estimation, distributed learning, online learning, differential privacy, group decision-making
conference: ; ; ccs: Security and privacyccs: Security and privacy Social network security and privacyccs: Theory of computationccs: Computing methodologies Multi-agent systems
1. Introduction
In many information environments, participating agents have privacy and security needs that prevent them from engaging in information exchange at their socially optimal levels (for optimum collective decision-making and social learning), e.g., when bound by client confidentiality, obligated by user privacy laws, or needing to safeguard their information sources. In other social and business settings, the participants may be reluctant to express unpopular opinions or openly deliberate on contentious or controversial issues. Even in the absence of such privacy, legal, security, or safety concerns, strategic and competitive considerations can lead to suboptimal information sharing at equilibrium. Differential privacy (DP) is a versatile framework for developing privacy-preserving algorithms, and it has been cardinal in large-scale and high-profile cases such as the U.S. decennial census (Garfinkel, 2022) and the tech industry (Erlingsson, 2014; Apple Differential Privacy Team, 2017). Earlier work also points to the potential of DP in alleviating misaligned incentives and reducing competitive inefficiencies in game-theoretic equilibria (Kearns et al., 2014).
Privacy is pivotal in many group decision scenarios. For example, consider intelligence operatives from allied countries who need to exchange information to reach a consensus on mission-critical objectives while safeguarding their information sources for the sustainability of their operations. This emphasis on privacy extends to public discourse in democracies, where the free exchange of information is crucial for effective deliberation but may never be realized if individuals face undue privacy risks. Privacy protections can enhance the quality of information aggregation for collective decision-making, particularly in environments where individuals harbor reasonable concerns about compromising their private information. For instance, when discussing topics such as substance abuse, mental health, gun violence, or abortion, privacy concerns may deter efficient participation to the extent that individuals’ opinions reflect their personal experiences. In other circumstances, individuals may face community retaliations or legal precautions that prevent them from expressing unpopular beliefs in public debates on contentious topics. In other settings, non-profit and for-profit organizations, while legally obligated to protect client data, can enhance service quality and operational profitability through responsible information sharing. Hospitals sharing health records can improve care quality and identify optimal treatment options for rare diseases, while banks collaborating on loan applicant records can enhance screening processes and mitigate the risk of loan defaults.
Our Contribution
In this work, we propose a differentially private [^1] In the sequel, we use DP to signify any of the phrases “differentially private”, “differentially privately”, and “differential privacy” with or without a hyphen between the words as appropriate., distributed, hypothesis-testing model for learning among a group of agents endowed with private signals and exchanging beliefs to infer a true state of the world or agree on a set of best alternatives. Our agents are wary of revealing their private signals or network neighborhoods to each other. The private signals are generated according to the true state, which is unknown but common to all agents. We propose non-Bayesian belief exchange models that protect private signals and network neighborhoods while achieving collective learning asymptotically.
We characterize the quality of learning subject to DP budget constraints and quantify the trade-off between learning rate and privacy: learning in privacy-critical environments can be sustained at a reasonable cost to communication complexity and accuracy. We show that non-Bayesian social learning models based on log-linear belief update (Gilardoni and Clayton, 1993; Rufo et al., 2012) are highly receptive to the privacy needs of the participating agents and can accommodate strict privacy requirements at a reasonable cost to social learners and their quality of group decision making.
DP constraints in our algorithms are satisfied by adding noise to belief updates at an appropriate level that protects individual signals or network neighborhoods, respecting the DP budget. In the online learning setting where agents have access to an infinite, intermittent stream of signals, the effect of noise vanishes asymptotically, but DP noising slows down the belief dynamics, requiring more communication rounds than the non-private baseline to achieve the same accuracy level. Our non-asymptotic (finite-time) analysis allows us to delineate the trade-offs between privacy budget, communication complexity, and error probability in the online learning case. When agents have access to only finitely many private signals, they aim to agree on a set of alternatives that best explains their collective observations using a distributed, maximum-likelihood estimation (distributed MLE). The crux of our analysis in this case lies in the fact that DP noising makes our distributed MLE algorithm non-deterministic. Hence, to achieve a desired accuracy level we need to repeat the distributed MLE algorithm in $K$ rounds. Our approach for DP distributed MLE opens a new design dimension to decide how best to aggregate the outcome of different rounds for a reliable MLE output. We propose two methods for aggregating the distributed MLE rounds that each offer distinct advantages in terms of communication complexity and their flexibility to control error probabilities.
Our first method is based on arithmetic and geometric averaging of beliefs across rounds. Specifically, it utilizes arithmetic averaging as a way to ensure that all good alternatives make their way to the identified set and offer a way to control the false positive rate (Type-I errors). On the other hand, geometric averaging offers a way to control missed detection rate (Type-II errors), ensuring that no bad alternatives are included in the MLE output.
Our second aggregation method is based on thresholding. Specifically, it keeps track of the frequencies that the beliefs in each round exceed a specific threshold. It then uses two thresholds to construct sets containing the states whose frequencies exceed either of the thresholds. This algorithm works analogously to the averaging algorithm and relies on the concentration of the frequencies to control Type-I or Type-II errors with either of the thresholds. The advantage of the thresholding method of aggregating over the averaging one is that it offers more flexibility to control Type-I and Type-II error rates individually; however, this comes at a cost to runtime and increased communication complexity.
Depending on the application, decision-makers can opt for their most suitable aggregators. For example, in dealing with rare and intractable diseases, health practitioners may be primarily concerned with identifying treatment options with the potential to improve patient outcomes and, thus, less concerned with false positives (inclusion of potentially ineffective treatment options in their identified alternatives). On the other hand, when screening loan applicants, financial institutions may be primarily concerned with avoiding loan defaults and may be open to missed loan opportunities to reduce the risk of bad investments. In the latter case, it would be desirable to have an aggregation method that ensures all bad states are filtered out with a lever to control missed detection rates.
In Section 2, we describe our networked information environment for group decision-making and provide non-private algorithms for distributed MLE and online learning that act as baselines for the remainder of the paper. In Section 3, we provide DP versions of these algorithms and offer performance analyses that delineate the trade-offs between communication complexity, privacy, and accuracy in each case. In Section 4, we provide a detailed literature review that contextualizes our results, followed by discussions of broader implications and conclusions in Section 5. Mathematical derivation and proof details are provided in the Appendix.
2. Preliminaries
2.1. Network Structure
We consider a set of $n$ agents, denoted by the set $[n]=\{1,\dots,n\}$, connected according to an undirected graph $\mathcal{G}([n],\mathcal{E})$ without self-loops. The neighborhood of agent $i$ is denoted by $\mathcal{N}_{i}$ and corresponds to all of the agents that agent $i$ can locally exchange information with. The graph corresponds to a doubly-stochastic adjacency matrix $A=\left[a_{ij}\right]_{i,j\in[n]}$ with weights $a_{ij}=0$ whenever $(i,j)\notin\mathcal{E}$ and weight $a_{ij}>0$ for all $(i,j)\in\mathcal{E}$, and moreover $\sum_{i\in[n]}a_{ij}=1$. The adjacency matrix $A$ is associated with $n$ eigenvalues $\lambda_{1}(A)\geq\lambda_{2}(A)\geq\dots\geq\lambda_{n}(A)$ and a set of bi-orthonormal eigenvectors $\{\bm{l}_{i},\bm{r}_{i}\}_{i\in[n]}$ satisfying $\|\bm{l}_{i}\|_{2}=\|\bm{r}_{i}\|_{2}=\bm{l}_{i}^{T}\bm{r}_{i}=1$ for all $i\in[n]$ and $\bm{l}_{i}^{T}\bm{r}_{j}=0$ for all $i\neq j$.
The weights $A$ can be chosen in many ways. One of the ways to choose the weights is to follow the Metropolis-Hastings algorithm and set $a_{ij}=1/\max\{\deg(j),\deg(i)\}$ for all $(i,j)\in\mathcal{E}$ and also $a_{ii}=1-\sum_{j\neq i}a_{ij}$.
2.2. Non-Private Algorithms
We consider two types of learning regimes: in the first setting, we are concerned with the calculation of the maximum likelihood states in a distributed way based on a set of initial observations, and in the second setting, we consider learning the MLE states in an online manner through repeatedly observing data streams. In the sequel, we present the two settings and the corresponding non-private algorithms, which serve as benchmarks for the DP regime.
Both algorithms are based on a log-linear update rule, in which an agent forms its belief by combining the likelihood of the observed signals, the belief of the agent from the previous iteration, and the beliefs of its neighbors. We also note that other update schemes, such as linear updates, are possible; however, they suffer from slower convergence (Kayaalp et al., 2022, Proposition 1).
Distributed MLE. In the distributed maximum likelihood estimation (MLE) estimation task, there is a set of likelihood maximizers $\Theta^{\star}\subset\Theta$ arbitrarily chosen by nature from a state space $\Theta$, which is publicly known to the agents. The agents aim to determine $\Theta^{\star}$ collectively by combining their own information and information from their local neighbors. Specifically, each agent has access to a dataset $\bm{S}_{i}$ of $n_{i}$ i.i.d. observations $\bm{S}_{i}=(\bm{s}_{1},\dots,\bm{s}_{n_{i}})$ where $\bm{s}_{1},\dots,\bm{s}_{n_{i}}\in\mathcal{S}$ obey a statistical model determined by the likelihood function
$\ell_{i}(\bm{S}_{i}|\hat{\theta})=\prod_{j=1}^{n_{i}}\ell_{i}(\bm{s}_{j}|\hat{\theta}).$
Our task is, given signals $\bm{S}_{i}$ for each agent $i\in[n]$, to find the set of likelihood maximizers, i.e.,
(1) $\Theta^{\star}=\operatorname*{argmax}_{\hat{\theta}\in\Theta}\Lambda(\hat{\theta}),\text{ where }\Lambda(\hat{\theta})=\sum_{i=1}^{n}\log\left(\ell_{i}(\bm{S}_{i}|\hat{\theta})\right).$
We define $\bar{\Theta}=\Theta\setminus\Theta^{\star}$ as the set of non-optimal states. We let $f^{\star}$ represent the density of MLE states, i.e., $f^{\star}=|\Theta^{\star}|/|\Theta|$.
In the non-private regime, the agents form beliefs $\bm{\mu}_{i,t}(\hat{\theta})$ over iteration $t$ and states and exchange their beliefs with their neighbors until they can detect the MLE. The work of (Rahimian and Jadbabaie, 2016b) gives an algorithm for belief exchange that provably converges to the true MLE after a finite number of steps (see Theorem 1 in their paper). The algorithm is given in Algorithm 1.
Algorithm 1 Non-Private Distributed MLE The agents begin by forming: $\bm{\gamma}_{i}(\tilde{\theta})=\prod_{j=1}^{n_{i}}\ell_{i}(\bm{s}_{j}|\tilde{\theta})$, and initializing their beliefs to ${\bm{\mu}}_{i,0}(\hat{\theta})$ $=$ $\bm{\gamma}_{i}(\hat{\theta})/\sum_{\tilde{\theta}\in\Theta}\bm{\gamma}_{i}(\tilde{\theta})$. In any future time period, the agents update their belief after communication with their neighboring agents, and according to the following update rule for any $\hat{\theta}$: (2) ${\bm{\mu}}_{i,t}(\hat{\theta})=\frac{{\bm{\mu}}^{1+a_{ii}}_{i,t-1}(\hat{\theta})\prod\limits_{j\in\mathcal{N}_{i}}{{\bm{\mu}}^{a_{ij}}_{j,t-1}(\hat{\theta})}}{\sum\limits_{\tilde{\theta}\in\Theta}{\bm{\mu}}_{i,t-1}^{1+a_{ii}}(\tilde{\theta})\prod\limits_{j\in\mathcal{N}_{i}}{{\bm{\mu}}^{a_{ij}}_{j,t-1}(\tilde{\theta})}}\cdot$
The convergence of this algorithm relies on studying the behavior of the log-belief ratio between two states $\hat{\theta}$ and $\check{\theta}$, whose dynamics are governed by the powers of the primitive matrix $A+I$ (see (Rahimian and Jadbabaie, 2016b, Theorem 1)) with ordered eigenvalues $\alpha_{i}=\lambda_{i}(A+I)$. Specifically, Theorem 1 of (Rahimian and Jadbabaie, 2016b) states that there is a dominant term due to the maximum eigenvalue $\alpha_{1}=2$ that dominates the log-belief ratios for large $t$, and $n-1$ terms that are exponentially decaying with rate $\alpha^{\star}_{n}=(1/2)\max\{\alpha_{2},|\alpha_{n}|\}<1$. The dominant term on each state depends on the difference of the log-likelihoods between the states $\hat{\theta}$ and $\check{\theta}$, namely $\Lambda(\hat{\theta},\check{\theta})=\Lambda(\hat{\theta})-\Lambda(\check{\theta})$. Subsequently, the log-belief ratio approaches $\infty$- ∞ as $t\to\infty$ whenever $\check{\theta}\in\Theta^{\star}$, and we can recover the MLEs.
[…]
Algorithm 2 Non-Private Distributed Online Learning Every time $t\in\mathbb{N}_{0}$, each agent forms the likelihood product of the signals that it has received at that iteration: $\bm{\gamma}_{i,t}(\hat{\theta})=\prod_{j=1}^{n_{i,t}}\ell_{i}(\bm{s}^{j}_{i,t}|\hat{\theta})$, if $n_{i,t}\geq 1$, and $\gamma_{i,t}(\hat{\theta})=1$ if $n_{i,t}=0$. The agent then updates its belief according to: (3) ${\bm{\mu}}_{i,t}(\hat{\theta})=\frac{\bm{\gamma}_{i,t}(\hat{\theta})\bm{\mu}^{a_{ii}}_{i,t-1}(\hat{\theta})\prod_{j\in\mathcal{N}_{i}}{{\bm{\mu}}^{a_{ij}}_{j,t-1}(\hat{\theta})}}{\sum\limits_{\tilde{\theta}\in\Theta}\bm{\gamma}_{i,t}(\tilde{\theta})\bm{\mu}^{a_{ii}}_{i,t-1}(\tilde{\theta})\prod_{j\in\mathcal{N}_{i}}{{\bm{\mu}}^{a_{ij}}_{j,t-1}(\tilde{\theta})}}\mathpunct{\raisebox{2.15277pt}{,}}$ initialized by: ${\bm{\mu}}_{i,0}(\hat{\theta})=\bm{\gamma}_{i,0}(\hat{\theta})/\sum_{\tilde{\theta}\in\Theta}\bm{\gamma}_{i,0}(\tilde{\theta})$.
For this algorithm, (Rahimian and Jadbabaie, 2016b) proves that it converges to learning the MLE states asymptotically. Their argument relies on the fact that the average log-belief ratio between a pair of states $\hat{\theta},\check{\theta}\in\Theta$ converges to the weighted sum of the KL divergences of all the agents.
2.3. Network Dynamics and Differential Privacy
Learning Dynamics. We consider two methods for differential privacy and refer to them as Signal/Data Differential Privacy (Signal/Data DP) and Network Differential Privacy (Network DP). Both algorithms are local in principle; the agents simply add multiplicative noise to their estimates to achieve a desired privacy guarantee. Roughly, Signal DP adds noise to protect the private signals of each agent, $\bm{S}_{i,t}$, and Network DP adds noise to protect its own private signals, $\bm{S}_{i,t}$, as well as the beliefs $\{\bm{\nu}_{j,t-1}(\hat{\theta})\}_{j\in\mathcal{N}_{i},\hat{\theta}\in\Theta}$ of its neighbors from round $t-1$. We assume that the non-private network dynamics evolve according to a multiplicative update rule:
(4) $\bm{\mu}_{i,t}(\hat{\theta})\propto F_{i,t}(\bm{\mu}_{i,t-1}(\hat{\theta}))\cdot G_{i,t}\left(\left\{\bm{\mu}_{j,t-1}(\hat{\theta})\right\}_{j\in\mathcal{N}_{i}}\right)\cdot H_{i,t}(\bm{S}_{i,t}|\hat{\theta}),$
for all $\hat{\theta}\in\Theta$, for each agent $i\in[n]$, and $t\geq 1$, where $F_{i,t}:\mathbb{R}\to(0,+\infty)$, $G_{i,t}:\mathbb{R}^{d_{i}}\to(0,+\infty)$, and $H_{i,t}:\mathcal{S}^{n_{i,t}}\to(0,+\infty)$ are functions determined by the learning algorithm, and correspond to the information from the agent’s own estimate, the information from the neighboring estimates, and the information from the agent’s private signal respectively.
To achieve differential privacy, each agent adds some amount of noise $\bm{d}_{i,t}(\hat{\theta})$ drawn from a distribution $\mathcal{D}_{i,t}(\hat{\theta})$ to their estimate and reports the noisy estimate to their neighbors. The noise is independent among the states and the agents. The agent can either aim to protect only their private signal — which we call Signal DP and denote by $\mathcal{M}^{S}$, or protect their network connections and their private signal — which we call Network DP and denote by $\mathcal{M}^{N}$. The noisy dynamics are
(5) $\bm{\nu}_{i,t}(\hat{\theta})\propto F_{i,t}(\bm{\nu}_{i,t-1}(\hat{\theta}))\cdot\underbrace{G_{i,t}\left(\left\{\bm{\nu}_{j,t-1}(\hat{\theta})\right\}_{j\in\mathcal{N}_{i}}\right)\cdot\overbrace{H_{i,t}(\bm{S}_{i,t}|\hat{\theta})}^{\text{{\color[rgb]{1,0,1}Signal DP $\mathcal{M}^{S}_{i,t}$}}}}_{\text{{\color[rgb]{1,.5,0}Network DP $\mathcal{M}^{N}_{i,T}$}}}\;\cdot\;e^{\bm{d}_{i,t}(\hat{\theta})}.$
Formally, the two types of mechanisms can also be written as
$\psi_{\mathcal{M}^{S}_{i,t}}(\bm{S}_{i,t})=\big{(}H_{i,t}(\bm{S}_{i,t}|\hat{\theta})\cdot e^{\bm{d}_{i,t}(\hat{\theta})}\big{)}_{\hat{\theta}\in\Theta}\text{ for {$\mathcal{M}^{S}_{i,t}$}},$
and
$\psi_{\mathcal{M}^{N}_{i,T}}\big{(}\bm{S}_{i,t},\left\{\bm{\nu}_{j,t-1}\right\}_{j\in\mathcal{N}_{i}}\big{)}=\big{(}H_{i,t}(\bm{S}_{i,t}|\hat{\theta})\cdot G_{i,t}\big{(}\{\bm{\nu}_{j,t-1}(\hat{\theta})\}_{j\in\mathcal{N}_{i}})\cdot e^{\bm{d}_{i,t}(\hat{\theta})}\big{)}_{\hat{\theta}\in\Theta}\text{ for {$\mathcal{M}^{N}_{i,T}$}}.$
Moreover, for a single sample $\bm{s}\in\mathcal{S}$, the definition of $({\varepsilon},\delta)$-differential privacy is given below
Definition 2.0 (Signal DP).
A mechanism is said to be $({\varepsilon},\delta)$-DP with respect to the signal if and only if for all $X\subseteq\mathbb{R}$ we have that
${\mathbb{P}[\psi_{\mathcal{M}^{S}_{i,t}}(\bm{s}_{i,t})\in X]}\leq e^{\varepsilon}{\mathbb{P}[\psi_{\mathcal{M}^{S}_{i,t}}(\bm{s}_{i,t}^{\prime})\in X]}+\delta,\text{ for all }\bm{s}_{i,t},\bm{s}_{i,t}^{\prime}\in\mathcal{S}_{i},\text{ s.t. }\left\|\bm{s}_{i,t}-\bm{s}_{i,t}^{\prime}\right\|_{1}\leq 1.$
Definition 2.0 (Network DP).
A mechanism is said to be $({\varepsilon},\delta)$-DP with respect to the network if and only if for all $X\subseteq\mathbb{R}$ we have that
${\mathbb{P}\big{[}\psi_{\mathcal{M}^{N}_{i,T}}\big{(}\bm{s}_{i,t},\left\{\bm{\nu}_{j,t-1}\right\}_{j\in\mathcal{N}_{i}}\big{)}\in X\big{]}}\leq e^{\varepsilon}{\mathbb{P}\big{[}\psi_{\mathcal{M}^{N}_{i,T}}\big{(}\bm{s}_{i,t}^{\prime},\left\{\bm{\nu}_{j,t-1}^{\prime}\right\}_{j\in\mathcal{N}_{i}}\big{)}\in X\big{]}}+\delta$
for all $\big{(}\bm{s}_{i,t},\left\{\bm{\nu}_{j,t-1}\right\}_{j\in\mathcal{N}_{i}}\big{)},\big{(}\bm{s}_{i,t}^{\prime},\big{\{}\bm{\nu}_{j,t-1}^{\prime}\big{\}}_{j\in\mathcal{N}_{i}}\big{)}\in\mathcal{S}_{i}\times\mathbb{R}^{\deg(i)}$, s.t.
$\big{\|}\big{(}\bm{s}_{i,t},\big{\{}\bm{\nu}_{j,t-1}\big{\}}_{j\in\mathcal{N}_{i}}\big{)}-\big{(}\bm{s}_{i,t}^{\prime},\big{\{}\bm{\nu}_{j,t-1}^{\prime}\big{\}}_{j\in\mathcal{N}_{i}}\big{)}\big{\|}_{1}\leq 1.$
Sensitivity. The global $\ell_{1}$-sensitivity of the sample log-likelihood of the $i$-th agent and state $\hat{\theta}\in\Theta$ is defined to be
$g_{i}(\hat{\theta})=\max_{\bm{s},\bm{s}^{\prime}\in\mathcal{S}:\|\bm{s}-\bm{s}^{\prime}\|_{1}\leq 1}\big{\|}\log\ell_{i}(\bm{s}|\hat{\theta})-\log\ell_{i}(\bm{s}^{\prime}|\hat{\theta})\big{\|}_{1}.$
The sensitivity of the joint log-likelihood is $n_{i,t}g_{i}(\hat{\theta})$ because each agent has access to $n_{i,t}$ i.i.d. observations at each point in time (in the general case) and each observation has point sensitivity $g_{i}(\hat{\theta})$.
In the context of Signal DP, the sensitivity of agent $i$ at iteration $t$ is defined to be
$\Delta_{i,t}=n_{i,t}\cdot\max_{\hat{\theta}\in\Theta}g_{i}(\hat{\theta}),$
as the worst-case sensitivity over the state space [^2] If we want to protect the number of signals, then we can take a max over all of the sensitivities $\Delta_{n,\Theta}=\max_{i\in[n],t\in\mathbb{N}}\Delta_{i,t}$. Moreover, even though we are not protecting the states, taking the maximum over $\Theta$ simplifies the theoretical analysis significantly (at the expense of adding more noise).. In the context of Network DP, the sensitivity of agent $i$ at iteration $t$ is instead defined to be
$\Delta_{i,t}=\max\left\{n_{i,t}\cdot\max_{\hat{\theta}\in\Theta}g_{i}(\hat{\theta}),1+\max_{j\neq i}a_{ij}\right\}.$
For both cases, we define the worst-case sensitivity to be $\Delta_{n,\Theta}=\max_{i\in[n],t\in\mathbb{N}}\Delta_{i,t}$.
$x$ Scalar
$\overline{\bm{x}}$ Vector
$\bm{x}$ Element of vector $\overline{\bm{x}}$
$\|\overline{\bm{x}}\|_{p}$ $\ell_{p}$-norm of $\overline{\bm{x}}$
$X$ Matrix
$n$ Number of agents
${\varepsilon}$ / $\delta$ Privacy budget and leakage probability
$\eta_{1}$ / $\eta_{2}$ Type-I/Type-II error probabilities
$A$ Communication matrix
$\Theta$ Set of states
$\Theta^{\star}$ / $\bar{\Theta}$ Set of MLE/non-MLE states
$\theta^{\star}$ / $\bar{\theta}$ MLE/non-MLE state
$f^{\star}$ Density of MLE states ($f^{\star}=|\Theta^{\star}|/|\Theta|$)
$\ell_{i}(\bm{s}|\hat{\theta})$ Likelihood of agent $i$ given signal $\bm{s}$ at state $\hat{\theta}$
$T$ / $K$ Number of iterations/rounds
$\bm{\gamma}_{i,t}(\hat{\theta})$ Joint Likelihood of agent $i$ given data $\bm{S}_{i,t}$ of size $n_{i,t}$ at state $\hat{\theta}$ in the non-private regime
$\bm{\sigma}_{i,t}(\hat{\theta})$ Joint Likelihood of agent $i$ given data $\bm{S}_{i,t}$ of size $n_{i,t}$ at state $\hat{\theta}$ in the private regime
$\bm{\mu}_{i,t}(\hat{\theta})$ Belief of agent $i$ at iteration $t$ on state $\hat{\theta}$ in the non-private regime
$\bm{\nu}_{i,k,t}(\hat{\theta})$ Belief of agent $i$ at iteration $t$ at round $k$ on state $\hat{\theta}$ in the private regime
$\Delta_{i,t}$ Sensitivity of agent $i$ at iteration $t$
$\bm{d}_{i,t}(\hat{\theta})$ / $\bm{d}_{i,k}(\hat{\theta})$ Noise variable for agent $i$, state $\hat{\theta}$, and iteration $t$/round $k$
$\mathcal{D}_{i,t}(\hat{\theta};{\varepsilon})$ / $\mathcal{D}_{i,k}(\hat{\theta};{\varepsilon})$ Noise distribution for agent $i$, state $\hat{\theta}$, and iteration $t$/round $k$
$\nu_{i}$ Expected value of number of signals that $i$ observes
$\xi_{i}$ Variance of number of signals that $i$ observes
$\beta_{n}^{\star}$ Second-largest eigenvalue modulus of $A$
$\alpha_{n}^{\star}$ Second-largest eigenvalue modulus of $(A+I)/2$
$\Gamma_{n,\Theta}$ Largest likelihood magnitude
$\Delta_{n,\Theta}$ Global sensitivity
$l_{n,\Theta}$ Minimum KL-divergence between any non-MLE state and an MLE state
$\Xi_{n}$ Sum of variances of numbers of signals accross all agents
$V_{n,\Theta}$ Sum of standard deviations of the noises over all agents
$Q_{n,\Theta}$ Maximum standard deviation of KL-divergence for any agent and any pair of states
Table 1. Notation table.
3. Differentially Private Algorithms
3.1. Performance Analysis Framework
In the sequel, we present our DP algorithms for distributed MLE and online learning of the true state from intermittent signals. We follow a similar approach for both distributed MLE and online learning by adding noise in the log domain. However, in the case of MLE, this noise makes the algorithm output random, so the algorithm needs to be repeated to achieve the desired accuracy level. In the online learning regime, the algorithm has access to a stream of noisy signals; therefore, the DP noise averages out over time. In the latter case, we only need to execute the algorithm once.
In addition to a fixed DP budget ${\varepsilon}>0$, our distributed MLE algorithms also respect fixed Type-I and Type-II error probabilities, denoted by $\eta_{1},\eta_{2}\in(0,1)$. The latter two parameters are because our randomized DP algorithms for distributed MLE can make two types of errors, which we control and bound in our analysis. For a distributed MLE algorithm $\mathcal{A}$ which returns an estimator $\hat{\Theta}^{\mathcal{A}}$, we define the Type-I error $\eta_{1}$ as the probability that $\left\{\Theta^{\star}\not\subseteq\hat{\Theta}^{\mathcal{A}}\right\}$, which can be controlled by bounding $\mathbb{P}\left[\Theta^{\star}\subseteq\hat{\Theta}^{\mathcal{A}}\right]\geq 1-\eta_{1}$. Similarly, we define the Type-II error $\eta_{2}$ to be the probability of the event $\left\{\Theta^{\star}\not\subseteq\hat{\Theta}^{\mathcal{A}}\right\}$, which can be controlled by ensuring that $\mathbb{P}\left[\Theta^{\star}\subseteq\hat{\Theta}^{\mathcal{A}}\right]\geq 1-\eta_{2}$. A Type-I error occurs as the algorithms more states than the true MLE set into its output. A Type-II error occurs as the algorithm filters out too many states, missing some MLE states in its output. In Section 1 we gave examples of cases in medical decision making and loan application screening where stakeholders may be interested in ensuring no MLE states are left out when dealing with intractable diseases (with a bound on false positive rate) or in ensuring that all bad states are filtered out to avoid loan defaults (with a lever to control missed detection rate). Similar scenarios arise in other application too, e.g., in business or military strategy one may prioritize removing false positive/type-I errors to avoid launching attacks without conclusive evidence, or otherwise, avoid missing the opportunity to capture high-value targets by removing missed detection/type-II errors.
To achieve guarantees on the DP budget and the two error probabilities, the algorithm needs to be run $T({\varepsilon},\eta)$ iterations and $K({\varepsilon},\eta)$ rounds, where $\eta$ here corresponds to either $\eta_{1}$ or $\eta_{2}$ (depending on the algorithm).
We define the communication complexity of an algorithm $\mathcal{A}$ to correspond to the total number of belief updates, which equals $K({\varepsilon},\eta)\cdot T({\varepsilon},\eta)$ in the distributed MLE case and $T({\varepsilon},\eta)$ in the online learning case. Table 2 shows the summary of our results for the two types of learning regimes for a fixed error probability $\eta$ (which either corresponds to the Type-I or the Type-II error) and fixed privacy budget ${\varepsilon}$.
Beyond the privacy budget ${\varepsilon}$, the error probability $\eta$, and the number of agents $n$, the results of Table 2 also depend on the statistical properties of the information environment and the network structure, as described below:
  • Regarding the private signal strcutures, the bounds depend on the largest $\log$-likelihood magnitude
    $\Gamma_{n,\Theta}=\max_{i\in[n],\hat{\theta}\in\Theta}|\log\bm{\gamma}_{i,t}(\hat{\theta})|,$
    the minimum divergence between any non-MLE state and an MLE state
    $l_{n,\Theta}=\min_{\bar{\theta}\in\bar{\Theta},\theta^{\star}\in\Theta^{\star}}\left|\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\bar{\theta})|\ell_{i}(\cdot|\theta^{\star})\right)\right|,$
    the global sensitivity $\Delta_{n,\Theta}=\max_{i\in[n],t\in\mathbb{N}}\Delta_{i,t}$, the sum of variances of the number of signals $\Xi_{n}=\sum_{i=1}^{n}\xi_{i}$, and the maximum deviation of the KL divergence [^3] By applying the main result of (Ghosal et al., 2000), one can relate $Q_{n,\Theta}$ to statistical distances.
    $Q_{n,\Theta}=\max_{i\in[n],\bar{\theta}\in\bar{\Theta},\theta^{\star}\in\Theta^{\star}}\sqrt{\mathbb{V}\left[\log\left(\frac{\ell_{i}(\cdot|\bar{\theta})}{\ell_{i}(\cdot|\theta^{\star})}\right)\right]}.$
  • Regarding the graph structure, our bounds depend on the second-largest eigenvalue modulus (SLEM) of $A$,
    $\beta^{\star}_{n}=\max\{\lambda_{2}(A),|\lambda_{n}(A)|\},$
    and the SLEM of $(A+I)/2$,
    $\alpha^{\star}_{n}=(1/2)\max\left\{\lambda_{2}(A+I),|\lambda_{n}(A+I)|\right\}.$
Algorithm With DP (ours) Without DP (Rahimian and Jadbabaie, 2016b)
Communication Complexity with respect to ${\varepsilon},\eta$
MLE (AM/GM) $\mathcal{O}_{|\Theta|,n}\left({\color[rgb]{0.16015625,0.5,0.7265625}{\log(1/\eta)(\log(1/\eta)+\log(1/{\varepsilon}))}}\right)$ $\mathcal{O}_{|\Theta|,n}\left({1}\right)$
MLE (Threshold) $\mathcal{O}_{|\Theta|,n}\left({\color[rgb]{0.16015625,0.5,0.7265625}{\frac{\log(1/\eta)}{\pi^{2}}(\log(1/\eta)+\log(1/{\varepsilon})+\log(1/q))}}\right)$ $\mathcal{O}_{|\Theta|,n}\left({1}\right)$
Online Learning $\mathcal{O}_{|\Theta|,n}\left({\color[rgb]{0.16015625,0.5,0.7265625}{\frac{1}{\eta}\sqrt{\frac{1}{{\varepsilon}}}}}\right)$ $\mathcal{O}_{|\Theta|,n}\left({1}\right)$
Communication Complexity with respect to $n$
MLE (AM/GM) $\mathcal{O}_{|\Theta|,{\varepsilon},\eta}\left(\max\left\{\log\left(\frac{n}{l_{n,\Theta}}\right),\frac{\log\left(n(\Gamma_{n,\Theta}+{\color[rgb]{0.16015625,0.5,0.7265625}{\Delta_{n,\Theta}}})\right)}{\log(1/\alpha^{\star}_{n})}\right\}\right)$ $\mathcal{O}_{|\Theta|}\left(\max\left\{\log\left(\frac{n}{l_{n,\Theta}}\right),\frac{\log\left({n}{\Gamma_{n,\Theta}}\right)}{\log(1/\alpha^{\star}_{n})}\right\}\right)$
[…]
Online Learning $\mathcal{O}_{|\Theta|,{\varepsilon},\eta}\left(\sqrt{\frac{n^{2}(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}})\left(Q_{n,\Theta}+{\color[rgb]{0.16015625,0.5,0.7265625}{\Delta_{n,\Theta}}}\right)}{l_{n,\Theta}(1-(\beta_{n}^{\star})^{2})}}\right)$ $\mathcal{O}_{|\Theta|}\left({\sqrt{\frac{n^{2}(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}})Q_{n,\Theta}}{l_{n,\Theta}(1-(\beta_{n}^{\star})^{2})}}}\right)$
Communication Complexity with respect to $|\Theta|$
[…]
Online Learning $Θ\mathcal{O}_{n,{\varepsilon},\eta}\left(\sqrt{\frac{(Q_{n,\Theta}+\Delta_{n,\Theta}){\color[rgb]{0.16015625,0.5,0.7265625}{|\Theta|}}}{l_{n,\Theta}}}\right)$ $\mathcal{O}_{n}\left({\sqrt{\frac{Q_{n,\Theta}}{l_{n,\Theta}}}}\right)$
Table 2. Summary of estimators for a fixed privacy budget ${\varepsilon}>0$ and target error $\eta\in(0,1)$. We use $\mathcal{O}_{a_{1},\dots,a_{N}}\left(\cdot\right)$ to denote the big-O notation where the constants are allowed to depend on $a_{1},\dots,a_{N}$. In our algorithms, we have set the threshold $\varrho$ to be a constant. The overhead of introducing privacy is outlined in blue. For compactness in notation, $\eta$ refers to $\max\{\eta_{1},\eta_{2}\}$, $\pi$ refers to $\min\{\pi_{2},\pi_{1}\}$, and $q$ refers to $\min\{1-q_{2},q_{1}\}$ (see Theorems 3.2 and 3.5). The AM/GM algorithms have matching communication complexities whenever $q=O(1)$ and $\pi=O(1/\sqrt{|\Theta|})$.
3.2. Distributed MLE
In Algorithm 3, we propose a DP version of Algorithm 1 for distributed MLE. The first idea behind the new algorithm is to introduce multiplicative noise subject to DP to the likelihood functions $\bm{\gamma}_{i}(\hat{\theta})$. Adding multiplicative noise – which corresponds to additive noise in the log domain – ensures privacy. Then, the agents communicate their beliefs over the states with their local neighbors and form estimates of the true MLE across time. However, introducing noise makes the algorithm randomized – compared to Algorithm 1, which is the deterministic counterpart – and therefore, the agents may misidentify the MLE with a non-trivial probability. To overcome this problem, we run the algorithm for $K$ independent rounds and combine these $K$ estimates to produce the final beliefs. By setting $K$ accordingly, we can reduce the algorithm’s error $\eta$, which we show in Theorem 3.1.
To produce the final estimates, we rely on two ways of aggregating the $K$ rounds of the algorithm: The first approach is to take the arithmetic mean of the $K$ rounds to produce the final belief, which we call the AM-Estimator. The benefit of AM-Estimator is that it can accurately w.h.p. identify all the MLE states, i.e., we can control its Type-II error probability: $\mathbb{P}\left[\Theta^{\star}\not\subseteq\hat{\Theta}^{\mathrm{AM}}_{i,T}\right]\leq\eta_{2}$. This estimator has high recall but has low precision because it can recover a superset of $\Theta^{\star}$. The second way to combine the $K$ rounds is to take the geometric mean, which we call the GM-Estimator. For the GM-Estimator we can control the Type-I error probability, $\mathbb{P}\left[\hat{\Theta}^{\mathrm{GM}}_{i,T}\not\subseteq\Theta^{\star}\right]\leq\eta_{1}$, to ensure it recovers a subset of $\Theta^{\star}$ w.h.p.
Algorithm 3 Private Distributed MLE (AM/GM) Inputs: Privacy budget ${\varepsilon}$, Error probabilities $\eta_{1},\eta_{2}$, Log-belief Threshold $\varrho>0$. Initialization: Set the number of iterations $T$ and the number of rounds $K$ as indicated by Theorem 3.2 (for the non-asymptotic case) or Theorem 3.1 (for the asymptotic case), and $\tau=1/(1+e^{\varrho})$. Procedure: The following is repeated in $K$ rounds indexed by $k\in[K]$. In each round $k$, agents begin by forming noisy log-likelihoods $\bm{\sigma}_{i,k}(\hat{\theta})=e^{\bm{d}_{i,k}(\hat{\theta})}\bm{\gamma}_{i}(\hat{\theta})$, where $\bm{\gamma}_{i}(\hat{\theta})=\prod_{j=1}^{{n}_{i}}\ell_{i}(\bm{s}^{j}_{i}|\tilde{\theta})$ and $\bm{d}_{i,k}(\hat{\theta})\sim\mathcal{D}_{i}(\hat{\theta};{\varepsilon})$ independently across agents $i\in[n]$ and states $\hat{\theta}\in\Theta$. The agents initialize their beliefs to ${\bm{\nu}}_{i,k,0}(\hat{\theta})$ $=$ $\bm{\sigma}_{i,k}(\hat{\theta})/\sum_{\tilde{\theta}\in\Theta}\bm{\sigma}_{i,k}(\tilde{\theta})$, and over the next $T$ time steps, they communicate with their neighbors and update their beliefs according to the following rule for any $\hat{\theta}$ and $t\in[T]$: (6) ${\bm{\nu}}_{i,k,t}(\hat{\theta})=\frac{{\bm{\nu}}^{1+a_{ii}}_{i,k,t-1}(\hat{\theta})\prod\limits_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,k,t-1}(\hat{\theta})}}{\sum\limits_{\tilde{\theta}\in\Theta}{\bm{\nu}}_{i,k,t-1}^{1+a_{ii}}(\tilde{\theta})\prod\limits_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,k,t-1}(\tilde{\theta})}}\cdot$ After $K$ rounds, the agents aggregate the outcome of the rounds for every $\hat{\theta}\in\Theta$ as follows: (7) ${\bm{\nu}}_{i,T}^{\mathrm{AM}}(\hat{\theta})=\frac{\sum_{k\in[K]}\bm{\nu}_{i,k,T}(\hat{\theta})}{K}\mathpunct{\raisebox{2.15277pt}{,}}\quad{\bm{\nu}}_{i,T}^{\mathrm{GM}}(\hat{\theta})=\frac{\prod_{k\in[K]}\bm{\nu}_{i,k,T}(\hat{\theta})^{1/K}}{\sum_{\tilde{\theta}\in\Theta}\prod_{k\in[K]}\bm{\nu}_{i,k,T}(\tilde{\theta})^{1/K}}\cdot$ Outputs: Return $\hat{\Theta}^{\mathrm{AM}}_{i,T}=\left\{\hat{\theta}\in\Theta:\bm{\nu}_{i,T}^{\mathrm{AM}}(\hat{\theta})\geq\tau\right\}\mathpunct{\raisebox{2.15277pt}{,}}\quad\hat{\Theta}^{\mathrm{GM}}_{i,T}=\left\{\hat{\theta}\in\Theta:\bm{\nu}_{i,T}^{\mathrm{GM}}(\hat{\theta})\geq\tau\right\}\cdot$
We first start with a result on the asymptotic behavior of Algorithm 3. Specifically, we show that by repeating the algorithm, we can recover the MLE with high probability. The reason for repeating the algorithm is that now the algorithm is randomized compared to the non-private version, and therefore, the algorithm if run for $K=1$ rounds, can fail with a positive probability since adding noise to the log domain can change the sign of the log-belief ratio between a non-MLE state $\bar{\theta}\in\bar{\Theta}$ and an MLE state $\theta^{\star}\in\Theta^{\star}$ with positive probability. We repeat the algorithm $K$ times and construct two estimators to aggregate the results to achieve this. We present the theorem:
Theorem 3.1 ().
For Algorithm 3, distributions $\mathcal{D}_{i}(\hat{\theta};{\varepsilon})=\mathcal{D}_{i}({\varepsilon})$ that satisfy ${\varepsilon}$-DP and do not depend on the state $\hat{\theta}$, we have that as $\varrho\to\infty$,
  • for $K\geq|\Theta^{\star}|\log(|\bar{\Theta}|/\eta_{1})$, we have that $\lim_{T\to\infty}\mathbb{P}\left[\hat{\Theta}^{\mathrm{GM}}_{i,T}\subseteq\Theta^{\star}\right]\geq 1-\eta_{1}$ for all $i\in[n]$.
  • for $K\geq|\bar{\Theta}|\log(|\Theta^{\star}|/\eta_{2})$, we have that $\lim_{T\to\infty}\mathbb{P}\left[\Theta^{\star}\subseteq\hat{\Theta}^{\mathrm{AM}}_{i,T}\right]\geq 1-\eta_{2}$ for all $i\in[n]$.
Proof Sketch. At a high level for the AM-Estimator, we show that the probability that an MLE state $\theta^{\star}\in\Theta^{\star}$ ends up with a non-zero belief at a given round $k\in[K]$ is at least $1/|\bar{\Theta}|$, and therefore we can show that the probability that the AM-Estimator gives a zero value to an MLE state is exponentially small, i.e., at most $e^{-K/|\bar{\Theta}|}$, since the algorithm is repeated $K$ times independently. Therefore, the probability that any MLE state is misclassified by the AM-Estimator is at most $|\Theta^{\star}|e^{-K/|\bar{\Theta}|}$ by the union bound. Making this probability at most $\eta_{2}$ yields the value of $K$. A similar argument can work for the GM-Estimator. ∎
The first observation is that when $f^{\star}>1/2$, namely the MLE states “dominate” $\Theta$, using the AM-Estimator requires fewer rounds than the GM-Estimator, and whenever $f^{\star}<1/2$, namely the non-MLE states “dominate” $\Theta$, running the GM-Estimator requires fewer rounds than the AM-Estimator, and when $f^{\star}=1/2$ both estimators need to be repeated the same number of times to get the guarantee.
As a second observation, when there is a majority of MLE states, recovering a superset of the MLE states with the AM-Estimator guarantees accuracy of at least $50\%$ due to a low Type-II error, and, on the other hand, when the non-MLE states are a majority, the using the GM-Estimator guarantees accuracy of at least $50\%$ in rejecting the non-MLE states due to a low Type-I error.
Finally, repeating Algorithm 3 for $K\geq|\Theta|\log(|\Theta|/\min\{\eta_{1},\eta_{2}\})$ iterations, one is able to assert that $\hat{\Theta}^{\mathrm{GM}}_{i,t}\subseteq\Theta^{\star}\subseteq\hat{\Theta}^{\mathrm{AM}}_{i,t}$ as $t\to\infty$ with probability $1-\eta_{1}-\eta_{2}$ without knowledge of $f^{\star}$.
It is interesting to point out that the above theorem holds regardless of the noise (as long as the noise distribution does not depend on the state for a given agent) and shows that randomizing the algorithm by adding noise deems the repetition of the algorithm essential for it being able to succeed w.h.p., contrary to the non-private version which is deterministic. Furthermore, choosing $\mathcal{D}_{i}({\varepsilon})=\mathrm{Lap}\left(\frac{K|\Theta|\Delta_{i}}{{\varepsilon}}\right)$ satisfies ${\varepsilon}$-DP and the assumptions of Theorem 3.1.
The previous analysis focused on the regime where $t\to\infty$. The next question that surfaces is studying the behavior of Algorithm 3 for finite $t$. As one would expect, the speed of convergence depends on the sum of standard deviations of the noise, i.e., $V_{n,\Theta}=\sum_{i=1}^{n}\sqrt{\mathbb{V}_{\bm{d}_{i}\sim\mathcal{D}_{i}({\varepsilon})}\left[\bm{d}_{i}\right]}$.
Following, we give a high-probability bound on recovering two sets $\hat{\Theta}^{\mathrm{GM}}_{i,T}$ and $\hat{\Theta}^{\mathrm{AM}}_{i,T}$, such that $\hat{\Theta}^{\mathrm{GM}}_{i,T}\subseteq\Theta^{\star}\subseteq\hat{\Theta}^{\mathrm{AM}}_{i,T}$:
Theorem 3.2 ().
The following hold for Algorithm 3:
  • For $T\geq\max\left\{\frac{\log\left(\frac{2\varrho n}{l_{n,\Theta}}\right)}{\log 2},\frac{\log\left(\frac{|\Theta|^{2}(n-1)(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\eta_{1}\varrho\sqrt{K}}\right)}{\log(1/\alpha^{\star}_{n})}\right\}$ and $K\geq|\Theta^{\star}|\log(|\bar{\Theta}|/\eta_{1})$, we have that $\mathbb{P}[\hat{\Theta}^{\mathrm{GM}}_{i,T}\subseteq\Theta^{\star}]\geq 1-2\eta_{1}$.
  • For $T\geq\max\left\{\frac{\log\left(\frac{2\varrho n}{l_{n,\Theta}}\right)}{\log 2},\frac{\log\left(\frac{|\Theta|^{2}(n-1)K(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\log(1/\eta_{2})\varrho}\right)}{\log(1/\alpha^{\star}_{n})}\right\}$ and $K\geq|\bar{\Theta}|\log(|\Theta^{\star}|/\eta_{2})$, we have that $\mathbb{P}[\Theta^{\star}\subseteq\hat{\Theta}^{\mathrm{AM}}_{i,T}]\geq 1-2\eta_{2}$.
  • The optimal distributions that minimize the convergence time $T$ for both estimators, are $\mathcal{D}_{i}^{\star}({\varepsilon})=\mathrm{Lap}\left(\frac{\Delta_{i}K|\Theta|}{{\varepsilon}}\right)$.
Proof Sketch. To derive the non-asymptotic result, we consider the log-belief ratio dynamics, which are linear with respect to the primitive matrix $A+I$. For each round $k\in[K]$, we can show that the log-belief ratio between any pair of states $\hat{\theta},\check{\theta}\in\Theta$ has two terms: The first term that depends on the principal eigenvalue $\alpha_{1}=2$ of the matrix $A+I$ and the noisy log-likelihood ratio $Z_{k}(\hat{\theta},\check{\theta})=\Lambda(\hat{\theta},\check{\theta})+\sum_{i=1}^{n}\bm{d}_{i,k}(\hat{\theta})-\sum_{i=1}^{n}\bm{d}_{i,k}(\check{\theta})$, and the second term decays with rate $\alpha^{\star}_{n}$ and depends on $\Gamma_{n,\Theta}$ and $V_{n,\Theta}$. To find the convergence time $T$ for each estimator, we first condition on the fact that $Z_{k}(\hat{\theta},\check{\theta})$ and $\Lambda(\hat{\theta},\check{\theta})$ have the same sign which yields a value for $K$ similar to Theorem 3.1, and then conditioned on this “good” event happening, we give a high-probability bound on the log-belief ratio between an MLE state $\theta^{\star}$ and a non-MLE state $\bar{\theta}$ which we make to be at least $\varrho$, and therefore the choice of threshold $\tau$ can return the desired sets ($\hat{\Theta}_{i,T}^{\mathrm{GM}}\subseteq\Theta^{\star}\subseteq\hat{\Theta}_{i,T}^{\mathrm{AM}}$) w.h.p. The distributions that optimize the runtime can be found by minimizing $V_{n,\Theta}$ subject to DP constraints and are proven to be Laplace with parameters $\Delta_{i}K|\Theta|/{\varepsilon}$ where $\Delta_{i}$ is the sensitivity of agent $i$, and ${\varepsilon}/(K|\Theta|)$ is the privacy budget that corresponds each round and each state due to the composition theorem. ∎
Remark about the threshold $\varrho$. To minimize the communication complexity, it suffices to pick the optimal threshold $\varrho^{\star}$ to be the one that makes the terms inside the maximum equal – and, hence, the maximum is minimized – which corresponds to the following two optimal thresholds:
$\varrho^{\star}_{\mathrm{AM}}$ $=\left(\frac{|\Theta|^{2}(n-1)K(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\log(1/\eta_{2})}\right)^{\frac{\log 2}{\log(2/\alpha_{n}^{\star})}}\left(\frac{l_{n,\Theta}}{2n}\right)^{\frac{\log(1/\alpha_{n}^{\star})}{\log(2/\alpha_{n}^{\star})}},$
$\varrho^{\star}_{\mathrm{GM}}$ $=\left(\frac{|\Theta|^{2}(n-1)(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\eta_{1}\sqrt{K}}\right)^{\frac{\log 2}{\log(2/\alpha_{n}^{\star})}}\left(\frac{l_{n,\Theta}}{2n}\right)^{\frac{\log(1/\alpha_{n}^{\star})}{\log(2/\alpha_{n}^{\star})}}.$
For simplicity of exposition, the results presented in Table 2 correspond to $\varrho$ being a constant number.
3.3. A Double-Threshold Algorithm for Recovering the MLE
In Algorithm 4 we provide a double-threshold algorithm to simultaneously control both type-I and type-II error probabilities, with more design flexibility that comes at an increased communication cost.
Algorithm 4 Private Distributed MLE (Double-Threshold Algorithm) Inputs: Privacy budget ${\varepsilon}$, Error probabilities $\eta_{1},\eta_{2}$, Log-belief Threshold $\varrho>0$. Initialization: Set the number of iterations $T$ and the number of rounds $K$, and the thresholds $\tau^{\mathrm{thres},2},\tau^{\mathrm{thres},1}$ as indicated by Theorem 3.5 (for the non-asymptotic case) or Theorem 3.3 (for the asymptotic case), and $\tau=1/(1+e^{\varrho})$. Procedure: The following is repeated in $K$ rounds indexed by $k\in[K]$. In each round $k$, the agents begin by forming the noisy log-likelihoods $\bm{\sigma}_{i,k}(\hat{\theta})=e^{\bm{d}_{i,k}(\hat{\theta})}\bm{\gamma}_{i}(\hat{\theta})$, where $\bm{\gamma}_{i}(\hat{\theta})=\prod_{j=1}^{{n}_{i}}\ell_{i}(\bm{s}^{j}_{i}|\tilde{\theta})$ and $\bm{d}_{i,k}(\hat{\theta})\sim\mathcal{D}_{i}(\hat{\theta};{\varepsilon})$ independently across agents $i\in[n]$ and states $\hat{\theta}\in\Theta$. The agents initialize their beliefs to ${\bm{\nu}}_{i,k,0}(\hat{\theta})$ $=$ $\bm{\sigma}_{i,k}(\hat{\theta})/\sum_{\tilde{\theta}\in\Theta}\bm{\sigma}_{i,k}(\tilde{\theta})$. Over the next $T$ time steps, the agents update their belief after communicating with their neighbors, and according to the following update rule for any $\hat{\theta}$ and all $t\in[T]$: (8) ${\bm{\nu}}_{i,k,t}(\hat{\theta})=\frac{{\bm{\nu}}^{1+a_{ii}}_{i,k,t-1}(\hat{\theta})\prod\limits_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,k,t-1}(\hat{\theta})}}{\sum\limits_{\tilde{\theta}\in\Theta}{\bm{\nu}}_{i,k,t-1}^{1+a_{ii}}(\tilde{\theta})\prod\limits_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,k,t-1}(\tilde{\theta})}}\cdot$ After $K$ rounds, each agent aggregates the outcome of the rounds for every $\hat{\theta}\in\Theta$ as follows: (9) $\bm{N}_{i,T}(\hat{\theta})$ $=\frac{1}{K}\sum_{k\in[K]}\mathds{1}\left\{\bm{\nu}_{i,k,T}(\hat{\theta})>\tau\right\}\cdot$ Outputs: Return $\hat{\Theta}^{\mathrm{thres},1}_{i,T}=\left\{\hat{\theta}\in\Theta:\bm{N}_{i,T}(\hat{\theta})\geq\tau^{\mathrm{thres},1}\right\}\mathpunct{\raisebox{2.15277pt}{,}}\quad\hat{\Theta}^{\mathrm{thres},2}_{i,T}=\left\{\hat{\theta}\in\Theta:\bm{N}_{i,T}(\hat{\theta})\geq\tau^{\mathrm{thres},2}\right\}\cdot$
The analysis of Theorem 3.1 shows that in a given round $k$, the probability an MLE state $\theta^{\star}\in\Theta$ ends up positive as $t\to\infty$ is at least $1/|\bar{\Theta}|$, so in expectation at least $K/|\bar{\Theta}|$ rounds will yield a positive belief. On the other hand, we know that for a non-MLE state $\bar{\theta}$, on expectation, at most $(1-1/|\Theta^{\star}|)K$ trials will come up heads in expectation. For brevity, we define $p_{1}=1-1/|\Theta^{\star}|$ and $p_{2}=1/|\bar{\Theta}|$.
Because $\bm{N}_{i,T}(\hat{\theta})$ is an average of independent indicator variables for all $\hat{\theta}\in\Theta$, the Chernoff bound indicates that it concentrates around its mean $\mathbb{E}\left[\bm{N}_{i,T}(\hat{\theta})\right]$. This prompts the development of the following simple algorithm, which resembles boosting algorithms such as AdaBoost (Freund and Schapire, 1997). The algorithm uses two thresholds $\tau^{\mathrm{thres},2},\tau^{\mathrm{thres},1}$ to estimate the true states as the states for which the number of times their beliefs have exceeded a threshold $\tau$ exceeds $\tau^{\mathrm{thres},2}$ and $\tau^{\mathrm{thres},1}$ respectively. The algorithm is summarized in Algorithm 4.
Our first result considers the asymptotic regime where the number of iterations $T$ goes to infinity. We present the theorem below:
Theorem 3.3 ().
For Algorithm 4, distributions $\mathcal{D}_{i}(\hat{\theta};{\varepsilon})=\mathcal{D}_{i}({\varepsilon})$ that are ${\varepsilon}$-DP and do not depend on the state $\hat{\theta}$, we have that as $\varrho\to\infty$,
  • for any $\pi_{1}>0$, $\tau^{\mathrm{thres},1}=(1+\pi_{1})p_{1}$, and $K\geq\frac{\log(|\bar{\Theta}|/\eta_{1})}{2\pi_{1}^{2}}$, we have that
    $\lim_{T\to\infty}\mathbb{P}\left[\hat{\Theta}^{\mathrm{thres},1}_{i,T}\subseteq\Theta^{\star}\right]\geq 1-\eta_{1},$
    for all $i\in[n]$.
  • for any $\pi_{2}>0$, $\tau^{\mathrm{thres},2}=(1-\pi_{2})p_{2}$, and $K\geq\frac{\log(|\Theta^{\star}|/\eta_{2})}{2\pi_{2}^{2}}$, we have that
    $\lim_{T\to\infty}\mathbb{P}\left[\Theta^{\star}\subseteq\hat{\Theta}^{\mathrm{thres},2}_{i,T}\right]\geq 1-\eta_{2},$
    for all $i\in[n]$.
Proof Sketch. The proof is very similar to Theorem 3.1, with the only difference that instead of relying on independent events, we use the Chernoff bound on $\bm{N}_{i,T}(\hat{\theta})$. ∎
Similarly to Theorem 3.1, the estimator $\hat{\Theta}^{1,\mathrm{thres}}_{i,T}$ has a low Type-I error and the estimator $\hat{\Theta}^{2,\mathrm{thres}}_{i,T}$ has a low Type-II error. If the agents do not have knowledge of $|\Theta^{\star}|$ they can use the thresholds ${\tau^{\mathrm{thres},1}}^{\prime}=(1+\pi_{1})p_{1}\geq\tau^{\mathrm{thres},1}$ and ${\tau^{\mathrm{thres},2}}^{\prime}=(1-\pi_{2})p_{2}\leq\tau^{\mathrm{thres},2}$ and obtain the same guarantee. Also, running the algorithm for $K\geq\max\left\{\frac{\log(|\Theta|/\eta_{1})}{2\pi_{1}^{2}}\mathpunct{\raisebox{2.15277pt}{,}}\frac{\log(|\Theta|/\eta_{2})}{2\pi_{2}^{2}}\right\}$ we can guarantee that as $T\to\infty$, we have $\Theta^{\mathrm{thres},2}_{i,T}\subseteq\Theta^{\star}\subseteq\Theta^{\mathrm{thres},1}_{i,T}$ with probability at least $1-\eta_{1}-\eta_{2}$.
Because $\bm{N}_{i,T}(\theta^{\star})$ for the MLE states are concentrated around a distribution with a mean that is at least $1/|\bar{\Theta}|$ and, similarly, $\bm{N}_{i,T}(\bar{\theta})$ for the non-MLE states are concentrated around their expectation which is at most $1-1/|\Theta^{\star}|$, one may rightly question whether it is possible to achieve perfect prediction of $\Theta^{\star}$ assuming the two modes are “sufficiently” well-separated. The answer to this is affirmative as long as $p_{2}>p_{1}$ and is given by the following Corollary:
Corollary 3.4 (Exact Recovery with a Single Threshold).
If the density of MLE states $f^{\star}$ satisfies
[…]
the two thresholds are set equal, i.e., $\tau^{\mathrm{thres},2}=\tau^{\mathrm{thres},1}$ which corresponds to $\hat{\Theta}^{\mathrm{thres},2}_{i,T}=\hat{\Theta}^{\mathrm{thres},1}_{i,T}$ and $\pi_{1}=(1-\pi_{2})\frac{p_{2}}{p_{1}}-p_{1}$ for some $\pi_{2}>0$, and $K\geq\max\left\{\frac{\log(|\Theta|/\eta_{1})}{2\pi_{1}^{2}},\frac{\log(|\Theta|/\eta_{2})}{2\pi_{2}^{2}}\right\}$ we have that
$\lim_{T\to\infty}\mathbb{P}\left[\hat{\Theta}^{\mathrm{thres},2}_{i,T}=\Theta^{\star}\right]\geq 1-\eta_{1}-\eta_{2}.$
Proof Sketch. The range of $f^{\star}$ is derived by solving the inequality $\frac{1}{(1-f^{\star})|\Theta|}>1-\frac{1}{f^{\star}|\Theta|}$, and the rest follows by applying Theorem 3.3. ∎
Furthermore, by performing an analysis similar to Theorem 3.2, we can derive a non-asymptotic result:
Theorem 3.5 ().
Let $q_{1},q_{2}\in(0,1)$. Then for Algorithm 4 the following hold:
  • For any $\pi_{1}>0$, $\tau^{\mathrm{thres},1}=(1+\pi_{1})q_{1}$, $T\geq\max\left\{\frac{\log\left(\frac{2\varrho n}{l_{n,\Theta}}\right)}{\log 2},\frac{\log\left(\frac{|\Theta|^{2}(n-1)(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2q_{1}\varrho}\right)}{\log(1/\alpha^{\star}_{n})}\right\}$, and $K\geq\frac{\log(|\bar{\Theta}|/\eta_{1})}{2\pi_{1}^{2}}$, we have that $\mathbb{P}\left[\hat{\Theta}^{\mathrm{thres},1}_{i,T}\subseteq\Theta^{\star}\right]\geq 1-\eta_{1}$.
  • For any $\pi_{2}>0$, $\tau^{\mathrm{thres},2}=(1-\pi_{2})q_{2}$, $T\geq\max\left\{\frac{\log\left(\frac{2\varrho n}{l_{n,\Theta}}\right)}{\log 2},\frac{\log\left(\frac{|\Theta|^{2}(n-1)(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2(1-q_{2})\varrho}\right)}{\log(1/\alpha^{\star}_{n})}\right\}$, and $K\geq\frac{\log(|\Theta^{\star}|/\eta_{2})}{2\pi_{2}^{2}}$, we have that $\mathbb{P}\left[\Theta^{\star}\subseteq\hat{\Theta}^{\mathrm{thres},2}_{i,T}\right]\geq 1-\eta_{2}$.
  • The optimal distributions that minimize the convergence time $T$ for both estimators, are $\mathcal{D}_{i}^{\star}({\varepsilon})=\mathrm{Lap}\left({\Delta_{i}K|\Theta|}/{{\varepsilon}}\right)$.
Proof Sketch. The proof utilizes the convergence results proved in Theorem 3.2, and the concentration of $\bm{N}_{i,T}(\hat{\theta})$ similarly to Theorem 3.3. ∎
Remarks. We can pick the log-belief thresholds as
$\varrho^{\star}_{\mathrm{thres},1}$ $=\left(\frac{|\Theta|^{2}(n-1)(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2q_{1}\sqrt{K}}\right)^{\frac{\log 2}{\log(2/\alpha_{n}^{\star})}}\left(\frac{l_{n,\Theta}}{2n}\right)^{\frac{\log(1/\alpha_{n}^{\star})}{\log(2/\alpha_{n}^{\star})}},$
$\varrho^{\star}_{\mathrm{thres},2}$ $=\left(\frac{|\Theta|^{2}(n-1)(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2(1-q_{2})\sqrt{K}}\right)^{\frac{\log 2}{\log(2/\alpha_{n}^{\star})}}\left(\frac{l_{n,\Theta}}{2n}\right)^{\frac{\log(1/\alpha_{n}^{\star})}{\log(2/\alpha_{n}^{\star})}},$
to minimize the number of iterations.
Further, we can show that we can achieve perfect recovery as long as $q_{2}>q_{1}$. Moreover, it i s easy to observe from the above that the communication complexity can be minimized by setting $1-q_{2}$ as close as possible to $q_{1}$. However, achieving perfect recovery of the MLE comes at a cost to the communication complexity, which has a polylogarithmic dependence on the inverse of the gap $|1-q_{2}-q_{1}|$.
3.4. Online Learning from Intermittent Streams
In Algorithm 5, we introduce the online learning version, where the agents have access to a data stream at every iteration. In this algorithm, the agents receive $n_{i,t}\geq 1$ signals at every iteration $t$, calculate the likelihood and the noisy version of it (in order to preserve DP) and then aggregate it with their self- and neighboring beliefs from iteration $t-1$. Here, the data are produced by a model with true parameters $\theta^{\circ}\in\Theta$, and the set of identified states is given as the singleton
$\theta^{\star}=\theta^{\circ}=\operatorname*{argmax}_{\hat{\theta}\in\Theta}\;\mathbb{E}_{\theta^{\circ}}\left[\Lambda(\hat{\theta})\right],$
for any true model $\theta^{\circ}\in\Theta$, assuming a collectively identifiable statistical model; collective in the sense that $\Lambda(\hat{\theta})$ in Equation 1 is defined as the sum of the individual log-likelihoods.
Algorithm 5 Private Online Learning Inputs: Privacy budget ${\varepsilon}$, Error probabilities $\eta$. Initialization: Set the number of iterations $T$ as in Theorem 3.7. Procedure: Every time $t\in\mathbb{N}_{0}$, each agent forms the likelihood product of the signals that it has received at that time period: $\bm{\gamma}_{i,t}(\hat{\theta})=\prod_{j=1}^{n_{i,t}}\ell_{i}(\bm{s}^{j}_{i,t}|\hat{\theta})$, if $n_{i,t}\geq 1$, and $\gamma_{i,t}(\hat{\theta})=1$ if $n_{i,t}=0$. The agent then draws a noise variable $\bm{d}_{i,t}(\hat{\theta})\sim\mathcal{D}_{i,t}(\hat{\theta};{\varepsilon})$ for all $i\in[n]$, $\hat{\theta}\in\Theta$ independently and forms $\bm{\sigma}_{i,t}(\hat{\theta})=e^{\bm{d}_{i,t}(\hat{\theta})}\bm{\gamma}_{i,t}(\hat{\theta})$ for all $\hat{\theta}\in\Theta$ (and set $\bm{\sigma}_{i,t}(\hat{\theta})=1$ if $n_{i,t}=0$). The agent then updates its belief for $T$ iterations according to: (11) ${\bm{\nu}}_{i,t}(\hat{\theta})=\frac{\bm{\sigma}_{i,t}(\hat{\theta})\bm{\nu}^{a_{ii}}_{i,t-1}(\hat{\theta})\prod_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,t-1}(\hat{\theta})}}{\sum\limits_{\tilde{\theta}\in\Theta}\bm{\sigma}_{i,t}(\tilde{\theta})\bm{\nu}^{a_{ii}}_{i,t-1}(\tilde{\theta})\prod_{j\in\mathcal{N}_{i}}{{\bm{\nu}}^{a_{ij}}_{j,t-1}(\tilde{\theta})}}\mathpunct{\raisebox{2.15277pt}{,}}$ initialized by: ${\bm{\nu}}_{i,0}(\hat{\theta})=\bm{\sigma}_{i,0}(\hat{\theta})/\sum_{\tilde{\theta}\in\Theta}\bm{\sigma}_{i,0}(\tilde{\theta})$. Outputs: After $T$ iterations, return $\hat{\theta}^{\mathrm{OL}}_{i,T}=\operatorname*{argmax}_{\hat{\theta}\in\Theta}\bm{\nu}_{i,t}(\hat{\theta})$.
For the asymptotic regime, the log-belief ratio between $\theta^{\star}$ and a non-MLE state $\bar{\theta}$ is going to converge to $(-1/n)\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\bar{\theta})|\ell_{i}(\cdot|\theta^{\star})\right)$ as $t\to\infty$ and the effect of DP noise is going to vanish as a consequence of the Césaro mean. The weak law of large numbers, and therefore the agents are going to learn the MLE states exponentially fast without the need to repeat the algorithm multiple times. For completeness, we state the asymptotic theorem and provide a proof sketch:
Theorem 3.6 ().
As $T\to\infty$, for any distributions $\mathcal{D}_{i,t}(\hat{\theta};{\varepsilon})=\mathcal{D}_{i,t}({\varepsilon})$ that satisfy ${\varepsilon}$-DP and do not depend on the state $\hat{\theta}$, Algorithm 5, we have that: If $\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\bar{\theta})|\ell_{i}(\cdot|\theta^{\star})\right)<0$ for all $\bar{\theta}\neq\theta^{\star}$, then $\lim_{T\to\infty}\bm{\nu}_{i,t}(\theta^{\star})=1$ and the agents learn $\theta^{\star}$ asymptotically fast with rate $-l_{n,\Theta}/n$.
Proof Sketch. We can show that the log-belief ratio $\overline{\bm{\psi}}_{t}(\hat{\theta},\check{\theta})$ between a pair of states $\hat{\theta}\in\Theta$ and $\check{\theta}\in\Theta$ obeys $\overline{\bm{\psi}}_{t}(\hat{\theta},\check{\theta})=\sum_{\tau=0}^{t}A^{\tau}\left(\overline{\bm{\lambda}}_{t-\tau}(\hat{\theta},\check{\theta})+\overline{\bm{d}}_{t-\tau}(\hat{\theta})-\overline{\bm{d}}_{t-\tau}(\check{\theta})\right)$ where $\overline{\bm{\lambda}}_{t-\tau}(\hat{\theta},\check{\theta})$ is the log-likelihood ratio between $\hat{\theta}$ and $\check{\theta}$ at iteration $t-\tau$ and $\overline{\bm{d}}_{t-\tau}(\hat{\theta}),\overline{\bm{d}}_{t-\tau}(\check{\theta})$ are the vectorized noise variables respectively. The Césaro mean together with the strong law of large numbers implies that $\lim_{t\to\infty}(1/t)\overline{\bm{\psi}}_{t}(\hat{\theta},\check{\theta})=(1/n)\mathds{1}^{T}\mathbb{E}\left[\overline{\bm{\lambda}}_{0}(\hat{\theta},\check{\theta})+\overline{\bm{d}}_{0}(\hat{\theta})-\overline{\bm{d}}_{0}(\check{\theta})\right]\mathds{1}=(1/n)\mathds{1}^{T}\mathbb{E}\left[\overline{\bm{\lambda}}_{0}(\hat{\theta},\check{\theta})\right]\mathds{1}$, as $\mathds{1}^{T}\mathbb{E}\left[\overline{\bm{d}}_{0}(\hat{\theta})-\overline{\bm{d}}_{0}(\check{\theta})\right]\mathds{1}=\mathds{1}^{T}\mathbb{E}\left[\overline{\bm{d}}_{0}(\hat{\theta})\right]\mathds{1}-\mathds{1}^{T}\mathbb{E}\left[\overline{\bm{d}}_{0}(\hat{\theta})\right]\mathds{1}=0$ because the noise does not depend on the states. Moreover, $\mathds{1}^{T}\mathbb{E}\left[\overline{\bm{\lambda}}_{0}(\hat{\theta},\check{\theta})\right]\mathds{1}=\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\hat{\theta})|\ell_{i}(\cdot|\check{\theta})\right)\mathds{1}$, and whenever $\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\bar{\theta})|\ell_{i}(\cdot|\theta^{\star})\right)<0$, the agents can learn the true state exponentially fast with rate $-l_{n,\Theta}/n$. ∎
The next theorem characterizes the non-asymptotic behavior of Algorithm 5:
Theorem 3.7 ().
Running Algorithm 5, for
$T=\mathcal{O}\bigg{(}\sqrt{\max\bigg{\{}\frac{n|\Theta|(nQ_{n,\Theta}+V_{n,\Theta})(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}/\eta})}{l_{n,\Theta}\eta(1-(\beta_{n}^{\star})^{2})},\sqrt{\frac{n\log(|\Theta|)}{l_{n,\Theta}}}\bigg{\}}}\bigg{)}$
iterations, yields $\mathbb{P}\left[\theta^{\star}=\hat{\theta}^{\mathrm{OL}}_{i,T}\right]\geq 1-\eta$ as long as $\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\bar{\theta})|\ell_{i}(\cdot|\theta^{\star})\right)<0$ for all $\bar{\theta}\neq\theta^{\star}$. Subsequently, the distributions that optimize the runtime are $\mathcal{D}_{i,t}^{\star}=\mathrm{Lap}\left({\Delta_{i,t}|\Theta|}/{{\varepsilon}}\right)$.
Proof Sketch. In this case, we can similarly prove that the dynamics in the log-belief ratio space are linear. The proof starts by showing that the deviation of $(1/t)\sum_{\tau=0}^{t-1}n_{i,\tau}$ from $\nu_{i}$ is no more than $\sqrt{\Xi_{n}/(\eta t)}$ with probability at least $1-\eta$. Subsequently, conditioned on this “good” event, we show that the distance between the average log-belief ratio $(1/t)\log\left(\bm{\nu}_{i,t}(\hat{\theta})/\bm{\nu}_{i,t}(\check{\theta})\right)$ between any pair of states converges to $(1/n)\sum_{i=1}^{n}\nu_{i}D_{KL}\left(\ell_{i}(\cdot|\hat{\theta})|\ell_{i}(\cdot|\check{\theta})\right)$ is inversely proportional to $\sqrt{t}$, $1-(\beta^{\star}_{n})^{2}$, and is proportional to the standard deviation of the KL divergence $Q_{n,\Theta}$ as well as $V_{n,\Theta}$. Similarly to Theorem 3.2, we threshold the log-belief ratio between $\theta^{\star}$ and any $\bar{\theta}\neq\theta^{\star}$ and obtain the convergence time $T$. The optimal distributions are derived similarly to Theorem 3.2. ∎
4. Related Work
In this section, we review the following collections of literature on privacy and group-decision making: (i) differential privacy, (ii) works at the intersection of group decision-making and social learning, (iii) privacy protections in social network contexts, and (iv) literature in distributed learning and estimation.
4.1. Differential Privacy
Data privacy has evolved to incorporate various concepts, including $K$-anonymity. In the modern definition, a mechanism is considered differentially private if it assigns similar records to the same value with equal likelihood. This definition has a significant implication: it ensures that the outcome of a statistical analysis remains consistent regardless of an individual’s participation in the social learning process. Numerous existing mechanisms, such as randomized response (Warner, 1965), the Laplace mechanism, and the Gaussian mechanism, can be demonstrated to adhere to differential privacy principles. Randomized response, for example, introduces random perturbations to binary responses, enabling the retrieval of the population means while allowing individual respondents to maintain plausible deniability. It serves as an example of adding noise to data.
When it comes to handling donated data for purposes like providing recommendations to public policy agencies or submitting online product reviews on e-commerce platforms, it is crucial to protect the privacy of data donors. This is because common anonymization techniques used in such scenarios are susceptible to different types of attacks, including identification risks (Barbaro et al., 2006), linkage and cross-referencing vulnerabilities (Sweeney, 1997, 2015), as well as statistical difference and re-identification attacks (Kumar et al., 2007). Therefore, safeguarding the privacy of donors’ information becomes an essential aspect of statistical disclosure control in these contexts. Our work contributes to the literature on DP by utilizing DP techniques to enable group decision-making in distributed environments.
4.2. Group Decision-Making and Social Learning
The problem of aggregating different individuals’ opinions and observations into a coherent collective option arises naturally in jury deliberations, expert committees, medical diagnoses, or board meetings where several stakeholders have to agree on a factual basis despite heterogeneity and uncertainty of their opinions and private information. A common way to resolve disagreements in such contexts is to engage in repeated peer interactions. Subsequently, a long line of literature going back to seminal works of (Aumann, 1976; Geanakoplos and Polemarchakis, 1982; DeGroot, 1974) investigates the formation and evolution of Bayesian and non-Bayesian beliefs towards consensus agreement. This problem has close parallels in distributed estimation (Borkar and Varaiya, 1982; Tsitsiklis and Athans, 1984), data fusion (McLaughlin et al., 2003), as well as consensus and coordination in distributed control (Jadbabaie et al., 2003; Mesbahi and Egerstedt, 2010). The formation, evolution, and aggregation of beliefs in distributed environments have attracted a lot of interest in statistics (Genest et al., 1984; Genest and Zidek, 1986; Garthwaite et al., 2005), engineering (Kayaalp et al., 2022; Moreau, 2005; Hatano and Mesbahi, 2005; Boyd et al., 2006; Xiao et al., 2005; Ren and Beard, 2005; Cortes et al., 2005; Rabbat et al., 2005; Tanner et al., 2007; Borkar and Varaiya, 1982; Tsitsiklis and Athans, 1984; Shahrampour et al., 2015), philosophy (Lehrer and Wagner, 1981; List and Puppe, 2009; Dietrich et al., 2016; Conradt et al., 2013), sociology (Hegselmann and Krause, 2002; Friedkin and Johnsen, 1990), machine learning (Stoica and Papadimitriou, 2021), and economics (Golub and Jackson, 2010; DeMarzo et al., 2003; Golub and Jackson, 2012; Jadbabaie et al., 2012; Banerjee et al., 2021), among others.
Of particular interest to us in this paper is the growing literature on non-Bayesian information aggregation and opinion pooling, going back to the classical work of DeGroot (DeGroot, 1974), who proposes linear opinion pools by averaging the latest beliefs of each agent with that of their neighbors. Several asymptotic properties and extensions of this model, including streams of private signals over time, are studied in the literature (Jadbabaie et al., 2012; Golub and Jackson, 2010; DeMarzo et al., 2003). To combine beliefs on a finite set of alternatives, we use geometric averaging and logarithmic opinion pools which also have a long history in Bayesian analysis and behavioral decision models (Gilardoni and Clayton, 1993; Rufo et al., 2012) and they can be justified under specific axioms (Molavi et al., 2018) or derived as a no-recall, behavioral update rule (Rahimian and Jadbabaie, 2015; Rahimian et al., 2015; Rahimian and Jadbabaie, 2016a). In this paper, we ask how one should limit the information requirement of a non-Bayesian update to guarantee differential privacy and still ensure consensus agreement and asymptotic learning for the agents.
4.3. Privacy in Social Network Contexts
With our formulation of DP for distributed learning environments, we can focus on privacy leaks through information flow on the network rather than the underlying network structure. More broadly, the fundamental relationship between information flow and privacy is well noted, cf., contextual integrity (Nissenbaum, 2009; Benthall and Cummings, 2022) or how social network contexts affect link formation and information sharing behaviors (Acemoglu et al., 2022, 2017). Notwithstanding, the privacy implications of information diffusion and algorithmic intervention on social networks are largely unexplored, except in a few studies (Rahimian et al., 2023b; Liell-Cock et al., 2020; Rahimian et al., 2023a). The existing work looks at this issue in specific, highly stylized scenarios. One study shows that it is extremely hard to hide the existence of a giant, connected component of infected nodes under an independent cascade model, and a simple inference attack can reveal the status of a good fraction of nodes (Rezaei et al., 2021). In another study, the authors propose a decaying pseudo-noise that is added locally to mask the influence structure against an external observer with access to the entire information flow under the Friedkin-Johnsen influence propagation model (Liell-Cock et al., 2020).
4.4. Literature in Distributed Learning and Estimation
An important distinction between estimating a continuous quantity in (Papachristou and Rahimian, 2023; Li et al., 2018; Kang et al., 2023) and choosing from a discrete set in this work is in the fact that DP randomization in the latter case induces a non-trivial failure probability that is bounded away from zero with increasing number of iterations. Hence, in our work, to satisfactorily control the quality of the group decision outcome, we need the agents to repeat their deliberations in multiple rounds and then aggregate the outcome of several rounds towards a collective belief. Different aggregation schemes are possible (e.g., based on arithmetic or geometric averaging of the beliefs and using different threshold rules to identify the selected alternatives), which lead to different false positive and missed detection rates and come with their own privacy guarantee and communication complexity requirements.
5. Discussion and Conclusions
First, for the AM/GM estimators, we conclude that introducing privacy in the distributed MLE task incurs a $\mathrm{polylog}\left(1/{\varepsilon},1/\eta\right)$ cost compared to the non-private benchmark with respect to ${\varepsilon}$ and $\eta$, as well as a $\mathrm{poly}\left(|\Theta|,\log(\Delta_{n,\Theta})\right)$ cost compared to the non-private benchmark with respect to $|\Theta|$. Second, in the double-threshold regime, the distributed MLE task incurs a $\mathrm{polylog}\left(1/{\varepsilon},1/\eta,1/q\right)$ cost with respect to ${\varepsilon}$ and $\eta$, as well as a $\mathrm{poly}\left(1/\pi,|\Theta|,\Delta_{n,\Theta},\log(1/q)\right)$ overhead with respect to $|\Theta|$.
We observe the following trade-off between the AM/GM algorithm (Algorithm 3) and the double-threshold algorithm (Algorithm 4): while the double-threshold algorithm has more free parameters and is therefore more flexible than the AM/GM algorithm, the added flexibility of the double-threshold algorithm comes at a cost to its communication complexity.
Moreover, in the online learning regime, we end up having a $\mathrm{poly}\left(1/{\varepsilon},1/\eta\right)$ dependency compared to the non-private benchmark with respect to ${\varepsilon}$ and $\eta$, and a $\mathrm{poly}\left(|\Theta|,\Delta_{n,\Theta}\right)$ dependency with respect to $|\Theta|$. This shows that achieving the same privacy and accuracy levels in the online case requires many more communication rounds than in the MLE case.
The above results prove the utility of log-linear opinion pools for group decision making when choosing from a set of alternatives in privacy-critical environments. Our algorithms can protect both private signals as well as the network neighborhoods (Signal and Network DP). Although signal DP can be achieved efficiently over time, protection of network neighborhood is limited to a one-time attack by an eavesdropper. This is due to the fact that beliefs are correlated over time and across network neighborhoods, whereas private signals are independent and remain protected at the ${\varepsilon}$-DP level by the post-processing immunity (Dwork and Roth, 2014, Proposition 2.1). Extending network protections over a finite ($N$) number of attacks is possible using the composition property (Dwork and Roth, 2014, Chapter 3) and by reducing the privacy budget to ${\varepsilon}/N$, which requires large noise levels and comes at a high cost to communication and accuracy.
In summary, our paper touches on topics at the intersections of group-decision making and privacy. While our work is purely theoretical in nature and does not pose – to the best of our knowledge – specific ethical challenges, the societal and policy impact of our work should be highlighted. Our work has applications for distributed hypothesis testing problems, such as, for example, testing drug treatments and loan screening, at which multiple entities want to estimate the true state of nature based on observations from their neighbors without exposing their own information and provide protocols for information fusion depending on the application.
Work on such sensitive applications may necessitate developing or adapting privacy regulations. Policymakers may need to consider balancing the benefits of collaborative decision-making with protecting individual privacy and intellectual property, especially given that our algorithms provide the correct answer with a high probability (but are still susceptible to negligible errors).
Our paper’s approach allows entities to benefit from each other’s private information without compromising data security. As a result, organizations can improve the quality of their services through enhanced decision-making without violating privacy obligations, positively impacting efficiency and cost savings. For instance, collaborative hypothesis testing in drug treatments might streamline research, reducing redundant efforts and costs.
Acknowledgements.
M.P. was partially supported by a LinkedIn Ph.D. Fellowship, an Onassis Fellowship (ID: F ZT 056-1/2023-2024), and grants from the A.G. Leventis Foundation and the Gerondelis Foundation. M.A.R. was partially supported by NSF SaTC-2318844. The authors would like to thank the seminar participants at Rutgers Business School, Jalaj Upadhyay, Jon Kleinberg, Kate Donahue, and Vasilis Charisopoulos for their valuable discussions and feedback.
Appendix A Distributed MLE
A.1. Log-Belief Ratio Notations
Let $\bm{\mu}_{i,t}(\hat{\theta})$ be the beliefs for the non-private system, and let $\bm{\nu}_{i,k,t}(\hat{\theta})$ be the private estimates for agent $i\in[n]$, round $k\in[K]$ and round $t\geq 0$. For the non-private algorithm (Algorithm 1) all pairs $\hat{\theta},\check{\theta}\in\Theta$ we let $\bm{\phi}_{i,t}(\hat{\theta},\check{\theta})=\log\left(\frac{\bm{\mu}_{i,t}(\hat{\theta})}{\bm{\mu}_{i,t}(\check{\theta})}\right)$ and $\bm{\lambda}_{i}(\hat{\theta},\check{\theta})=\log\left(\frac{\bm{\gamma}_{i}(\hat{\theta})}{\bm{\gamma}_{i}(\check{\theta})}\right)$ be the log belief ratios and the log-likelihood ratios respectively for agent $i\in[n]$ round $t\geq 0$. For the private algorithm (Algorithm 3) we let $\bm{\psi}_{i,k,t}(\hat{\theta},\check{\theta})=\log\left(\frac{\bm{\nu}_{i,k,t}(\hat{\theta})}{\bm{\nu}_{i,k,t}(\check{\theta})}\right),\bm{\zeta}_{i}(\hat{\theta},\check{\theta})=\log\left(\frac{\bm{\sigma}_{i,k}(\hat{\theta})}{\bm{\sigma}_{i,k}(\check{\theta})}\right),$ and $\bm{\kappa}_{i,k}(\hat{\theta},\check{\theta})=\bm{d}_{i,k}(\hat{\theta})-\bm{d}_{i,k}(\check{\theta})$ be the private log-belief ratio, log-likelihood ratio and noise difference ratios respectively for agent $i\in[n]$, round $k\in[K]$ and round $t\geq 0$. We also let the vectorized versions $\overline{\bm{\phi}}_{t}(\hat{\theta},\check{\theta}),\overline{\bm{\psi}}_{k,t}(\hat{\theta},\check{\theta}),\overline{\bm{\lambda}}(\hat{\theta},\check{\theta}),\overline{\bm{\zeta}}_{k}(\hat{\theta},\check{\theta})$ respectively, where the vectorization is over the agents $i\in[n]$. We define $\Lambda(\hat{\theta},\check{\theta})=\mathds{1}^{T}\overline{\bm{\lambda}}(\hat{\theta},\check{\theta})$, $Z_{k}(\hat{\theta},\check{\theta})=\mathds{1}^{T}\overline{\bm{\zeta}}_{k}(\hat{\theta},\check{\theta})=\mathds{1}^{T}\left(\overline{\bm{\lambda}}(\hat{\theta},\check{\theta})+\overline{\bm{\kappa}}_{k}(\hat{\theta},\check{\theta})\right)$.
[…]
A.2. Auxiliary Lemmas
To prove the results we need the following auxiliary lemma:
Lemma A.1 ().
Let $X_{1},\dots,X_{n}$ be i.i.d. random variables drawn from a distribution $\mathcal{D}$. Then, for every $i\in[n]$, we have that $\mathbb{P}[X_{i}\geq\max_{j\neq i}X_{j}]=1/n$.
Proof.
Let $E_{i}=\{X_{i}\geq\max_{j\neq i}X_{j}\}$. Note that because $X_{1},\dots,X_{n}$ are i.i.d. $\mathbb{P}[E_{1}]=\mathbb{P}[E_{2}]=\dots=\mathbb{P}[E_{n}]$. Moreover, note that since the maximum is unique, we have that $E_{1},\dots,E_{n}$ are a partition of the sample space. Therefore, we get that $\sum_{i\in[n]}\mathbb{P}[E_{i}]=1$, and subsequently $\mathbb{P}[E_{i}]=1/n$. ∎
A.3. Proof of Theorem 3.1
Proof.
Similarly to (Rahimian and Jadbabaie, 2016b, Theorem 1), we have that for any $k\in[K]$
$\left\|\overline{\bm{\psi}}_{k,t}(\hat{\theta},\check{\theta})-\frac{2^{t}}{n}Z_{k}(\hat{\theta},\check{\theta})\mathds{1}\right\|_{2}$ $=\left\|\sum_{i=2}^{n}\left(\frac{\alpha_{i}}{2}\right)^{t}\overline{\bm{l}}_{i}\overline{\bm{r}}_{i}^{T}\overline{\bm{\zeta}}_{k}(\hat{\theta},\check{\theta})\right\|_{2}$
$\leq\sum_{i=2}^{n}\left|\frac{\alpha_{i}}{2}\right|^{t}\|\overline{\bm{l}}_{i}\|_{2}\left|\overline{\bm{r}}_{i}^{T}\overline{\bm{\zeta}}_{k}(\hat{\theta},\check{\theta})\right|$
$\leq\sum_{i=2}^{n}\left|\frac{\alpha_{i}}{2}\right|^{t}\left\|\overline{\bm{\zeta}}_{k}(\hat{\theta},\check{\theta})\right\|_{2}$
$\leq\sum_{i=2}^{n}\left|\frac{\alpha_{i}}{2}\right|^{t}\left(\left\|\overline{\bm{\lambda}}(\hat{\theta},\check{\theta})\right\|_{2}+\left\|\overline{\bm{\kappa}}_{k}(\hat{\theta},\check{\theta})\right\|_{2}\right).$
Taking expectations and applying Jensen’s inequality, we can bound the above as
(12) $\mathbb{E}\left[\left\|\overline{\bm{\psi}}_{k,t}(\hat{\theta},\check{\theta})-\frac{2^{t}}{n}Z_{k}(\hat{\theta},\check{\theta})\mathds{1}\right\|_{2}\right]\leq 2(n-1)(\alpha^{\star}_{n})^{t}\left[n\Gamma_{n,\Theta}+\sum_{i=1}^{n}\sqrt{\mathbb{V}\left[\bm{d}_{i,k}(\hat{\theta})\right]}\right],$
where $\Gamma_{n,\Theta}=\max_{i\in[n],\hat{\theta}\in\Theta}|\log\bm{\gamma}_{i}(\hat{\theta})|$. Note that the right-hand side in Equation 12 goes to 00 as $t\to\infty$, which implies that (by Markov’s Inequality) $\overline{\bm{\psi}}_{k,t}(\hat{\theta},\check{\theta})\overset{L_{2}}{\to}\frac{2^{t}}{n}Z_{k}(\hat{\theta},\check{\theta})\mathds{1}$.
Let $\theta^{\star}\in\Theta^{\star}$ and $\bar{\theta}\in\bar{\Theta}$. Based on our definition of $\Lambda(\cdot,\cdot)$ we should have $\Lambda(\bar{\theta},\theta^{\star})<0$ and the corresponding non-private algorithm would have $\phi_{i,t}(\bar{\theta},\theta^{\star})\to-\infty$ for every $\bar{\theta}\in\bar{\Theta}$ and $\theta^{\star}\in\Theta^{\star}$. However, when noise is introduced, there could be mistakes introduced in the log-belief ratios, even if $\Lambda(\bar{\theta},\theta^{\star})<0$ we have that $Z_{k}(\bar{\theta},\theta^{\star})\geq 0$ which will imply that $\bm{\psi}_{i,k,t}(\bar{\theta},\theta^{\star})\not\to-\infty$.
AM-Estimator. For all $\hat{\theta}\in\Theta$ we let $\bm{Y}_{j}(\hat{\theta})=\sum_{i=1}^{n}\bm{d}_{i,k}(\hat{\theta})$. We focus on a single $\theta^{\star}\in\Theta^{\star}$. It is easy to show that because the noise is independent on the agents, and i.i.d. on the states for a given agent $i$, then $\bm{Y}_{j}(\bar{\theta})$ and $\bm{Y}_{j}(\theta^{\star})$ are also i.i.d. for all $\bar{\theta}\in\bar{\Theta}$, and therefore we have that (see Lemma A.1)
$\lim_{T\to\infty}\mathbb{P}\left[\bm{\nu}_{i,k,T}(\theta^{\star})>0\right]$ $\geq\mathbb{P}\left[\bigcap_{\bar{\theta}\in\bar{\Theta}}\left\{\bm{Y}_{k}(\bar{\theta})\leq\bm{Y}_{k}(\theta^{\star})\right\}\right]=\frac{1}{|\bar{\Theta}|}$
Therefore, for the AM estimator we have that
(13) $\lim_{T\to\infty}\mathbb{P}\left[\bm{\nu}^{\mathrm{AM}}_{i,T}(\theta^{\star})=0\right]=\lim_{T\to\infty}\mathbb{P}\left[\bigcap_{k\in[K]}\left\{\bm{\nu}_{i,k,T}(\theta^{\star})=0\right\}\right]\leq\left(1-\frac{1}{|\bar{\Theta}|}\right)^{K}\leq e^{-\frac{K}{|\bar{\Theta}|}}$
and the failure probability of the AM estimator can be calculated by applying the union bound as
$\lim_{T\to\infty}\mathbb{P}\left[\Theta^{\star}\not\subseteq\hat{\Theta}^{\mathrm{AM}}_{i,T}\right]$ $=\lim_{T\to\infty}\mathbb{P}\left[\exists\theta^{\star}\in\Theta^{\star}:\bm{\nu}^{\mathrm{AM}}_{i,T}(\theta^{\star})=0\right]$
$\leq\lim_{T\to\infty}\sum_{\theta^{\star}\in\Theta^{\star}}\mathbb{P}\left[\bm{\nu}^{\mathrm{AM}}_{i,T}(\theta^{\star})=0\right]$
$\leq|\Theta^{\star}|e^{-\frac{K}{|\bar{\Theta}|}}.$
Setting $K=|\bar{\Theta}|\log(|\Theta^{\star}|/\eta_{2})$, the Type-II error is at most $\eta_{2}$.
GM-Estimator. Similarly, for $\bm{\nu}_{i,T}^{\mathrm{GM}}$ we have that
$\lim_{T\to\infty}\mathbb{P}\left[\bm{\nu}_{i,k,T}(\bar{\theta})=0\right]\geq\mathbb{P}\left[\bigcap_{\theta^{\star}\in\Theta^{\star}}\left\{\bm{Y}_{k}(\bar{\theta})\leq\bm{Y}_{k}(\theta^{\star})\right\}\right]=\frac{1}{|\Theta^{\star}|}$
for all $\bar{\theta}\in\bar{\Theta}$ and subsequently
$\lim_{T\to\infty}\mathbb{P}\left[\bm{\nu}_{i,T}^{\mathrm{GM}}(\bar{\theta})>0\right]=\lim_{T\to\infty}\mathbb{P}\left[\bigcap_{k\in[K]}\left\{\bm{\nu}_{i,k,T}(\bar{\theta})>0\right\}\right]\leq\left(1-\frac{1}{|\Theta^{\star}|}\right)^{K}\leq e^{-\frac{K}{|\Theta^{\star}|}},$
and
$\lim_{T\to\infty}\mathbb{P}\left[\hat{\Theta}^{\mathrm{GM}}_{i,T}\not\subseteq\Theta^{\star}\right]$ $=\lim_{T\to\infty}\mathbb{P}\left[\hat{\Theta}^{\mathrm{GM}}_{i,T}\cap\bar{\Theta}\neq\emptyset\right]$
$=\lim_{T\to\infty}\mathbb{P}\left[\exists\bar{\theta}\in\bar{\Theta}:\bm{\nu}^{\mathrm{GM}}_{i,T}(\bar{\theta})>0\right]$
$\leq\lim_{T\to\infty}\sum_{\bar{\theta}\in\bar{\Theta}}\mathbb{P}\left[\bm{\nu}^{\mathrm{GM}}_{i,T}(\bar{\theta})>0\right]$
$\leq|\bar{\Theta}|e^{-\frac{K}{|\Theta^{\star}|}}.$
Therefore, picking $K=|\Theta^{\star}|\log(|\bar{\Theta}|/\eta_{1})$ makes the Type-I error at most $\eta_{1}$.
A.4. Proof of Theorem 3.2
Proof.
GM-Estimator. We let $\bm{Y}_{k}(\hat{\theta})=\sum_{i\in[n]}\bm{d}_{i,k}(\hat{\theta})$.
We define the event $F=\left\{\exists k\in[K],\bar{\theta}\in\bar{\Theta},\theta^{\star}\in\Theta^{\star}:Z^{\mathrm{GM}}(\bar{\theta},\theta^{\star})>\Lambda(\bar{\theta},\theta^{\star})\right\}$. We have that
$\mathbb{P}[F]$ $\stackrel{{\scriptstyle\text{(i)}}}{{\leq}}\sum_{\bar{\theta}\in\bar{\Theta}}\mathbb{P}\left[\exists k\in[K],\theta^{\star}\in\Theta^{\star}:Z_{k}(\bar{\theta},\theta^{\star})>\Lambda(\bar{\theta},\theta^{\star})\right]$
=$\stackrel{{\scriptstyle\text{(ii)}}}{{=}}\sum_{\bar{\theta}\in\bar{\Theta}}\left(1-\mathbb{P}\left[\forall\theta^{\star}\in\Theta^{\star}:Z_{1}(\bar{\theta},\theta^{\star})\leq\Lambda(\bar{\theta},\theta^{\star})\right]\right)^{K}$
$\stackrel{{\scriptstyle\text{(iii)}}}{{\leq}}|\bar{\Theta}|\left(1-\mathbb{P}\left[\bigcap_{\bar{\theta}\in\bar{\Theta}}\left\{\bm{Y}_{1}(\bar{\theta})\leq\bm{Y}_{1}(\theta^{\star})\right\}\right]\right)^{K}$
$\stackrel{{\scriptstyle\text{(iv)}}}{{\leq}}|\bar{\Theta}|\left(1-\frac{1}{|\Theta^{\star}|}\right)^{K}$
$\stackrel{{\scriptstyle\text{(v)}}}{{\leq}}|\bar{\Theta}|e^{-K/|\Theta^{\star}|},$
where the above follows (i) from the application of the union bound, (ii) using the fact that the $K$ rounds are independent, (iii) the fact that $\{Z_{1}(\bar{\theta},\theta^{\star})\leq\Lambda(\bar{\theta},\theta^{\star})\}\supseteq\bigcap_{\bar{\theta}\in\bar{\Theta}}\left\{\bm{Y}_{k}(\bar{\theta})\leq\bm{Y}_{k}(\theta^{\star})\right\}$ and that $\{\bm{Y}_{k}(\bar{\theta})\}_{\bar{\theta}\in\bar{\Theta}}$ and $\bm{Y}_{k}(\theta^{\star})$ are i.i.d., (iv) Lemma A.1, and (v) $1+x\leq e^{x}$ for all $x\in\mathbb{R}$.
For each round $k$ of the algorithm and each pair of states $\hat{\theta},\check{\theta}\in\Theta$, the following holds for the log-belief ratio of the geometrically averaged estimates:
$\bm{\psi}_{i,t}^{\mathrm{GM}}(\hat{\theta},\check{\theta})=\frac{1}{K}\sum_{k\in[K]}\bm{\psi}_{i,k,t}(\hat{\theta},\check{\theta}).$
We let $Z^{\mathrm{GM}}(\hat{\theta},\check{\theta})=\frac{1}{K}\sum_{k=1}^{K}Z_{k}(\hat{\theta},\check{\theta})$. By the triangle inequality and Theorem 3.1
(14) $\mathbb{E}\left[\left\|\overline{\bm{\psi}}_{t}(\hat{\theta},\check{\theta})-\frac{2^{t}}{n}Z^{\mathrm{GM}}(\hat{\theta},\check{\theta})\mathds{1}\right\|_{2}\right]$ $\leq\frac{1}{\sqrt{K}}\sum_{k=1}^{K}\mathbb{E}\left[\left\|\overline{\bm{\psi}}_{k,t}(\hat{\theta},\check{\theta})-\frac{2^{t}}{n}Z_{k}(\hat{\theta},\check{\theta})\mathds{1}\right\|_{2}\right].$
For every round $k\in[K]$,
$\left\|\overline{\bm{\psi}}_{k,t}(\hat{\theta},\check{\theta})-\frac{2^{t}}{n}Z_{k}(\hat{\theta},\check{\theta})\mathds{1}\right\|_{2}$ $=\left\|\sum_{i=2}^{n}\left(\frac{\alpha_{i}}{2}\right)^{t}\overline{\bm{l}}_{i}\overline{\bm{r}}_{i}^{T}\overline{\bm{\zeta}}_{k}(\hat{\theta},\check{\theta})\right\|_{2}$
$\leq\sum_{i=2}^{n}\left|\frac{\alpha_{i}}{2}\right|^{t}\|\overline{\bm{l}}_{i}\|_{2}\left|\overline{\bm{r}}_{i}^{T}\overline{\bm{\zeta}}_{k}(\hat{\theta},\check{\theta})\right|$
$\leq\sum_{i=2}^{n}\left|\frac{\alpha_{i}}{2}\right|^{t}\left\|\overline{\bm{\zeta}}_{k}(\hat{\theta},\check{\theta})\right\|_{2}$
$\leq\sum_{i=2}^{n}\left|\frac{\alpha_{i}}{2}\right|^{t}\left(\left\|\overline{\bm{\lambda}}(\hat{\theta},\check{\theta})\right\|_{2}+\left\|\overline{\bm{\kappa}}_{k}(\hat{\theta},\check{\theta})\right\|_{2}\right).$
Taking expectations and applying Jensen’s inequality, we can bound the above as
$\mathbb{E}\left[\left\|\overline{\bm{\psi}}_{k,t}(\hat{\theta},\check{\theta})-\frac{2^{t}}{n}Z_{k}(\hat{\theta},\check{\theta})\mathds{1}\right\|_{2}\right]\leq 2(n-1)(\alpha^{\star}_{n})^{t}\left[n\Gamma_{n,\Theta}+\sum_{i=1}^{n}\sqrt{\mathbb{V}\left[\bm{d}_{i,k}(\hat{\theta})\right]}\right],$
Therefore Equation 14 becomes
(15) $\mathbb{E}\left[\left\|\overline{\bm{\psi}}_{t}(\hat{\theta},\check{\theta})-\frac{2^{t}}{n}Z^{\mathrm{GM}}(\hat{\theta},\check{\theta})\mathds{1}\right\|_{2}\right]$ $\leq\frac{2(n-1)(\alpha^{\star}_{n})^{t}[n\Gamma_{n,\Theta}+V_{n,\Theta}]}{\sqrt{K}}.$
By Markov’s inequality and Equation 14, we get that for every $z>0$
$\mathbb{P}\left[\left\|\overline{\bm{\psi}}_{t}^{\mathrm{GM}}(\hat{\theta},\check{\theta})-\frac{2^{t}}{n}Z^{\mathrm{GM}}(\hat{\theta},\check{\theta})\mathds{1}\right\|_{2}>z\right]\leq\frac{2(n-1)(\alpha^{\star}_{n})^{t}[n\Gamma_{n,\Theta}+V_{n,\Theta}]}{z\sqrt{K}}.$
We let the RHS be equal to $\eta_{1}/(|\Theta^{\star}||\bar{\Theta}|$, which corresponds to letting $z=|\bar{\Theta}||\Theta^{\star}|\frac{2(n-1)(\alpha^{\star}_{n})^{t}[n\Gamma_{n,\Theta}+V_{n,\Theta}]}{\eta_{1}\sqrt{K}}$. By applying a union bound over $\bar{\Theta}$ and $\Theta^{\star}$ we have that for every $\bar{\theta}\in\bar{\Theta}$ and $\theta^{\star}\in\Theta^{\star}$, with probability at least $1-\eta_{1}$,
(16) $\bm{\psi}_{i,t}^{\mathrm{GM}}(\bar{\theta},\theta^{\star})$ $\leq\frac{2^{t}}{n}Z^{\mathrm{GM}}(\bar{\theta},\theta^{\star})+|\bar{\Theta}||\Theta^{\star}|\frac{2(n-1)(\alpha^{\star}_{n})^{t}[n\Gamma_{n,\Theta}+V_{n,\Theta}]}{\eta_{1}\sqrt{K}}.$
Conditioning on $F^{c}$, the above becomes
$\bm{\psi}_{i,t}^{\mathrm{GM}}(\bar{\theta},\theta^{\star})$ $\leq\frac{2^{t}}{n}\Lambda(\bar{\theta},\theta^{\star})+|\bar{\Theta}||\Theta^{\star}|\frac{2(n-1)(\alpha^{\star}_{n})^{t}[n\Gamma_{n,\Theta}+V_{n,\Theta}]}{\eta_{1}\sqrt{K}}$
(17) $\leq-\frac{2^{t}}{n}l_{n,\Theta}+|\Theta|^{2}\frac{(n-1)(\alpha^{\star}_{n})^{t}[n\Gamma_{n,\Theta}+V_{n,\Theta}]}{2\eta_{1}\sqrt{K}}.$
To make Equation 17 at most $-\varrho$ for some $\varrho>0$, we need to set
$t\geq\max\left\{\frac{\log\left(\frac{2\varrho n}{l_{n,\Theta}}\right)}{\log 2},\frac{\log\left(\frac{|\Theta|^{2}(n-1)(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\eta_{1}\varrho\sqrt{K}}\right)}{\log(1/\alpha^{\star}_{n})}\right\}=T.$
The log-belief ratio threshold implies that for all $\theta^{\star}\in\Theta^{\star},\bar{\theta}\in\bar{\Theta}$ we have that $\bm{\nu}_{i,k,T}(\theta^{\star})\geq e^{\varrho}\bm{\nu}_{i,k,T}(\bar{\theta})$. To determine the belief threshold $\tau$, note that
11 11 $=\sum_{\hat{\theta}\in\hat{\Theta}^{\mathrm{GM}}_{i,T}}\bm{\nu}_{i,k,T}(\hat{\theta})+\sum_{\hat{\theta}\notin\hat{\Theta}^{\mathrm{GM}}_{i,T}}\bm{\nu}_{i,k,T}(\hat{\theta})\geq(1+e^{\varrho})\max_{\hat{\theta}\notin\Theta^{\mathrm{GM}}_{i,T}}\bm{\nu}_{i,k,T}(\hat{\theta})$
$\implies\max_{\hat{\theta}\notin\Theta^{\mathrm{GM}}_{i,T}}\bm{\nu}_{i,k,T}(\hat{\theta})\leq\frac{1}{1+e^{\varrho}}.$
Moreover, we can prove that $\min_{\hat{\theta}\in\hat{\Theta}^{\mathrm{GM}}_{i,T}}\bm{\nu}_{i,k,T}(\hat{\theta})\geq\frac{1}{1+e^{-\varrho}}\geq\frac{1}{1+e^{\varrho}}$ since $\varrho>0$ which shows that any value in $[1/(1+e^{\varrho}),1/(1+e^{-\varrho})]$ is a valid threshold. This yields that $\mathbb{P}[\hat{\Theta}^{\mathrm{GM}}_{i,T}\subseteq\Theta^{\star}|F^{c}]=\mathbb{P}[\hat{\Theta}^{\mathrm{GM}}_{i,T}\cap\bar{\Theta}=\emptyset|F^{c}]=\mathbb{P}[\exists\bar{\theta}\in\bar{\Theta}:\bm{\nu}_{i,T}^{\mathrm{GM}}(\bar{\theta})>\tau|F^{c}]\geq 1-\eta_{1}$. Subsequently, by letting $K=|\Theta^{\star}|\log(|\bar{\Theta}|/\eta_{1})$, we get that
$\mathbb{P}[\hat{\Theta}^{\mathrm{GM}}_{i,T}\subseteq\Theta^{\star}]\geq\mathbb{P}[\hat{\Theta}^{\mathrm{GM}}_{i,T}\subseteq\Theta^{\star}|F^{c}]\mathbb{P}[F^{c}]\geq(1-\eta_{1})^{2}\geq 1-2\eta_{1}.$
AM-Estimator. We let $E=\left\{\exists k\in[K],\bar{\theta}\in\bar{\Theta},\theta^{\star}\in\Theta^{\star}:Z_{k}(\theta^{\star},\bar{\theta})<\Lambda(\theta^{\star},\bar{\theta})\right\}$. Using similar arguments to Theorem 3.2, we can deduce that $\mathbb{P}[E]\leq|\Theta^{\star}|e^{-K/|\bar{\Theta}|}$. By setting $K=|\bar{\Theta}|\log(|\Theta^{\star}|/\eta_{2})$, we make $\mathbb{P}[E]\leq\eta_{2}$.
Conditioned on $E^{c}$ and by applying Markov’s inequality similarly to Theorem 3.2, we have that for all $\theta^{\star}\in\Theta^{\star},\bar{\theta}\in\bar{\Theta}$ and run $k\in[K]$
(18) $\bm{\psi}_{i,k,t}(\theta^{\star},\bar{\theta})$ $\geq\frac{2^{t}}{n}Z_{k}(\theta^{\star},\bar{\theta})-|\Theta^{\star}||\bar{\Theta}|\frac{2(n-1)(\alpha^{\star})^{t}(n\Gamma_{n,\Theta}+V_{n,\Theta})}{\eta_{2}^{\prime}}$
$\geq\frac{2^{t}}{n}\Lambda(\theta^{\star},\bar{\theta})-|\Theta|^{2}\frac{(n-1)(\alpha^{\star}_{n})^{t}(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\eta_{2}^{\prime}}$
$\geq\frac{2^{t}}{n}l_{n,\Theta}-|\Theta|^{2}\frac{2(n-1)(\alpha^{\star}_{n})^{t}(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\eta_{2}^{\prime}},$
with probability $1-\eta_{2}^{\prime}$ for some $\eta_{2}^{\prime}\in(0,1)$. To make the above at least $\varrho$ for some $\varrho>0$ it suffices to pick
$t\geq\max\left\{\frac{\log\left(\frac{2\varrho n}{l_{n,\Theta}}\right)}{\log 2},\frac{\log\left(\frac{|\Theta|^{2}(n-1)(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\eta_{2}^{\prime}\varrho}\right)}{\log(1/\alpha^{\star}_{n})}\right\}=T.$
The log-belief ratio threshold implies that for all $\theta^{\star}\in\Theta^{\star},\bar{\theta}\in\bar{\Theta}$ we have that $\bm{\nu}_{i,k,T}(\theta^{\star})\geq e^{\varrho}\bm{\nu}_{i,k,T}(\bar{\theta})$. Similarly, any value in $[1/(1+e^{\varrho}),1/(1+e^{-\varrho})]$ is a valid threshold, we can show that
$\mathbb{P}[\Theta^{\star}\subseteq\hat{\Theta}^{\mathrm{AM}}_{i,T}|E^{c}]$ $=\mathbb{P}\left[\forall\theta^{\star}\in\Theta^{\star}:\bm{\nu}^{\mathrm{AM}}_{i,T}(\theta^{\star})>{\tau}\bigg{|}E^{c}\right]$
$\geq\mathbb{P}\left[\bigcup_{k\in[K]}\left\{\forall\theta^{\star}\in\Theta^{\star}:\bm{\nu}_{i,k,T}(\theta^{\star})>\tau\right\}\bigg{|}E^{c}\right]$
$=1-\mathbb{P}\left[\bigcap_{k\in[K]}\left\{\exists\theta^{\star}\in\Theta^{\star}:\bm{\nu}_{i,k,T}(\theta^{\star})<\tau\right\}\bigg{|}E^{c}\right]$
$=1-\left(\mathbb{P}\left[\exists\theta^{\star}\in\Theta^{\star}:\bm{\nu}_{i,k,T}(\theta^{\star})<\tau\bigg{|}E^{c}\right]\right)^{K}$
$=1-\left(1-\mathbb{P}\left[\forall\theta^{\star}\in\Theta^{\star}:\bm{\nu}_{i,k,T}(\theta^{\star})>\tau\bigg{|}E^{c}\right]\right)^{K}$
$=1-(1-\eta_{2}^{\prime})^{K}$
$\geq 1-e^{-\eta_{2}^{\prime}K}.$
We set $\eta_{2}^{\prime}=\log(1/\eta_{2})/K$ and have that
$\mathbb{P}[\Theta^{\star}\subseteq\hat{\Theta}^{\mathrm{AM}}_{i,T}]\geq\mathbb{P}[E^{c}]\mathbb{P}[\Theta^{\star}\subseteq\hat{\Theta}^{\mathrm{AM}}_{i,T}|E^{c}]\geq(1-\eta_{2})^{2}\geq 1-2\eta_{2}.$
These finally yield
$T=\max\left\{\frac{\log\left(\frac{2\varrho n}{l_{n,\Theta}}\right)}{\log 2},\frac{\log\left(\frac{|\Theta|^{2}(n-1)K(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2\log(1/\eta_{2})\varrho}\right)}{\log(1/\alpha^{\star}_{n})}\right\}.$
Privacy (for both estimators). Minimizing the convergence time $T$ corresponds to minimizing $V_{n,\Theta}$. Since $V_{n,\Theta}$ is separable over the agents’ distributions, it suffices to solve the problem of minimizing the variance of each noise variable independently.
If an adversary can eavesdrop only once and has access to any round $k\in[K]$, and we have $|\Theta|$ states, by the composition theorem, the budget ${\varepsilon}$ should be divided by $K|\Theta|$. Thus, the problem of finding the optimal distribution $\mathcal{D}_{i}({\varepsilon})$ (with PDF say $p_{\mathcal{D}_{i}({\varepsilon})}(x)\in\Delta(\mathbb{R})$) corresponds to the following optimization problem studied in (Koufogiannis et al., 2015; Papachristou and Rahimian, 2023), for all $k\in[K]$:
(19) $\min_{p_{\mathcal{D}_{i}({\varepsilon})}(\cdot)\in\Delta(\mathbb{R})}$ $\quad\mathbb{V}_{\bm{d}_{i,k}\sim\mathcal{D}_{i}({\varepsilon})}\left[\bm{d}_{i,k}\right]$
s.t. $\quad\mathcal{D}_{i}({\varepsilon})\text{ is $\frac{{\varepsilon}}{K|\Theta|}$-DP and $\mathbb{E}_{\bm{d}_{i,k}\sim\mathcal{D}_{i}({\varepsilon})}\left[\bm{d}_{i,k}\right]=0$}.$
The optimal solution of this problem is given in (Koufogiannis et al., 2015; Papachristou and Rahimian, 2023). For the Signal DP problem, this corresponds to selecting $\mathcal{D}_{i}^{\star}({\varepsilon})=\mathrm{Lap}\left(\frac{\Delta_{i}K|\Theta|}{{\varepsilon}}\right)$ where $\Delta_{i}$ is the sensitivity of the log-likelihood of the $i$-th agent.
A.5. Proof of Theorem 3.3
Proof.
For brevity, we define $p_{2}=1/|\bar{\Theta}|$ and $p_{1}=1-1/|\Theta^{\star}|$.
Type-II Estimator ($\hat{\Theta}^{\mathrm{thres},2}_{i,T}$). To determine the asymptotic Type-II error probability $\eta_{2}$ of $\hat{\Theta}^{\mathrm{thres},2}_{i,T}$, we assume that the threshold takes the form $\tau^{\mathrm{thres},2}=(1-\pi_{2})p_{2}$. Then
(20) $\lim_{T\to\infty}\mathbb{P}\left[\Theta^{\star}\not\subseteq\hat{\Theta}^{\mathrm{thres},2}_{i,T}\right]$ =$\stackrel{{\scriptstyle\text{(i)}}}{{=}}\lim_{T\to\infty}\mathbb{P}\left[\exists\theta^{\star}\in\Theta^{\star}:\theta^{\star}\notin\Theta^{\mathrm{thres},2}_{i,T}\right]$
$\stackrel{{\scriptstyle\text{(ii)}}}{{\leq}}\lim_{T\to\infty}\sum_{\theta^{\star}\in\Theta^{\star}}\mathbb{P}\left[\theta^{\star}\notin\Theta^{\mathrm{thres},2}_{i,T}\right]$
$\stackrel{{\scriptstyle\text{(iii)}}}{{\leq}}\lim_{T\to\infty}\sum_{\theta^{\star}\in\Theta^{\star}}\mathbb{P}\left[\bm{N}_{i,T}(\theta^{\star})\leq\tau^{\mathrm{thres},2}\right]$
$\stackrel{{\scriptstyle\text{(iii)}}}{{\leq}}\lim_{T\to\infty}\sum_{\theta^{\star}\in\Theta^{\star}}\mathbb{P}\left[\bm{N}_{i,T}(\theta^{\star})\leq(1-\pi_{2})p_{2}\right]$
$\stackrel{{\scriptstyle\text{(iv)}}}{{\leq}}\lim_{T\to\infty}\sum_{\theta^{\star}\in\Theta^{\star}}\mathbb{P}\left[\bm{N}_{i,T}(\theta^{\star})\leq(1-\pi_{2})\mathbb{E}\left[\bm{N}_{i,T}(\theta^{\star})\right]\right]$
$\stackrel{{\scriptstyle\text{(v)}}}{{\leq}}|\Theta^{\star}|e^{-2K\pi_{2}^{2}},$
where the result is derived by applying (i) the definition of $\hat{\Theta}^{\mathrm{thres},2}_{i,T}$, (ii) union bound, (iii) the definition of the threshold $\tau^{\mathrm{thres},2}=(1-\pi_{2})p_{2}$, (iv) the fact that $\mathbb{E}\left[\bm{N}_{i,T}(\theta^{\star})\right]\geq p_{2}$ for all $\theta^{\star}\in\Theta^{\star}$ as $T\to\infty$, and (v) the Chernoff bound on $\bm{N}_{i,T}(\hat{\theta})$. Therefore, to make the above less than $\eta_{2}$, it suffices to choose $K=\frac{\log(|\bar{\Theta}|/\eta_{2})}{2\pi_{2}^{2}}$.
Type-I Estimator ($\hat{\Theta}^{\mathrm{thres},1}_{i,T}$). To determine the asymptotic Type-II error $\eta_{1}$ of $\hat{\Theta}^{\mathrm{thres},1}_{i,T}$, we assume that the threshold takes the form $\tau^{\mathrm{thres},1}=(1+\pi_{1})p_{1}$. Then
(21) $\lim_{T\to\infty}\mathbb{P}\left[\hat{\Theta}^{\mathrm{thres},1}_{i,T}\not\subseteq\Theta^{\star}\right]$ =$\stackrel{{\scriptstyle\text{(i)}}}{{=}}\lim_{T\to\infty}\mathbb{P}\left[\exists\bar{\theta}\in\bar{\Theta}:\bar{\theta}\in\hat{\Theta}^{\mathrm{thres},1}_{i,T}\right]$
$\stackrel{{\scriptstyle\text{(ii)}}}{{\leq}}\lim_{T\to\infty}\sum_{\bar{\theta}\in\bar{\Theta}}\mathbb{P}\left[\bar{\theta}\in\hat{\Theta}^{\mathrm{thres},1}_{i,T}\right]$
$\stackrel{{\scriptstyle\text{(iii)}}}{{\leq}}\lim_{T\to\infty}\sum_{\bar{\theta}\in\bar{\Theta}}\mathbb{P}\left[\bm{N}_{i,T}(\bar{\theta})\geq\tau^{\mathrm{thres},1}\right]$
$\leq\lim_{T\to\infty}\sum_{\bar{\theta}\in\bar{\Theta}}\mathbb{P}\left[\bm{N}_{i,T}(\bar{\theta})\geq(1+\pi_{1})p_{1}\right]$
$\stackrel{{\scriptstyle\text{(iv)}}}{{\leq}}\lim_{T\to\infty}\sum_{\bar{\theta}\in\bar{\Theta}}\mathbb{P}\left[\bm{N}_{i,T}(\bar{\theta})\geq(1+\pi_{1})\mathbb{E}\left[N_{i,T}(\bar{\theta})\right]\right]$
$\stackrel{{\scriptstyle\text{(v)}}}{{\leq}}|\bar{\Theta}|e^{-2K\pi_{1}^{2}},$
where the result is derived by applying (i) the definition of $\hat{\Theta}^{\mathrm{thres},1}_{i,T}$, (ii) union bound, (iii) the definition of the threshold $\tau^{\mathrm{thres},1}=(1+\pi_{1})p_{1}$, (iv) the fact that $\mathbb{E}\left[\bm{N}_{i,T}(\bar{\theta})\right]\leq p_{1}$ for all $\bar{\theta}\in\bar{\Theta}$ as $T\to\infty$, and the Chernoff bound on $\bm{N}_{i,T}(\bar{\theta})$. Therefore, to make the above $\eta_{1}$, it suffices to choose $K=\frac{\log(|\bar{\Theta}|/\eta_{1})}{2\pi_{1}^{2}}$.
A.6. Proof of Theorem 3.5
Type-II estimator. From the analysis of Theorem 3.2 (see Equation 18), setting
$t\geq\max\left\{\frac{\log\left(\frac{2\varrho n}{l_{n,\Theta}}\right)}{\log 2},\frac{\log\left(\frac{(n-1)|\Theta|^{2}(n\Gamma_{n,\Theta}+V_{n,\Theta})}{2(1-q_{2})\varrho}\right)}{\log(1/\alpha^{\star}_{n})}\right\}=T$
guarantees that $\mathbb{E}\left[\bm{N}_{i,T}(\theta^{\star}\right]=\mathbb{P}\left[\bm{\nu}_{i,k,T}(\theta^{\star})>\tau\right]\geq q_{2}$ for all $\theta^{\star}\in\Theta^{\star}$. Then by following the analysis similar to Theorem 3.5 by utilizing the union bound and the Chernoff bound, we can prove that $\mathbb{P}[\Theta^{\star}\not\subseteq\hat{\Theta}^{\mathrm{thres},2}_{i,T}]\leq|\Theta^{\star}|e^{-2K\pi_{2}^{2}}$. Subsequently, setting $K\geq\frac{\log(|\Theta^{\star}|/\eta_{2})}{2\pi_{2}^{2}}$ makes the error at most $\eta_{2}$.
Type-I estimator. From the analysis of Theorem 3.2 (see Equation 17), setting
$t\geq\max\left\{\frac{\log\left(\frac{2\varrho n}{l_{n,\Theta}}\right)}{\log 2},\frac{\log\left(\frac{(n-1)|\Theta|^{2}(n\Gamma_{n,\Theta}+V_{n,\Theta})}{q_{1}\varrho}\right)}{\log(1/\alpha^{\star}_{n})}\right\}=T$
guarantees that $\mathbb{E}\left[\bm{N}_{i,T}(\bar{\theta})\right]=\mathbb{P}\left[\bm{\nu}_{i,k,T}(\bar{\theta})\leq\tau\right]\leq q_{1}$ for all $\bar{\theta}\in\bar{\Theta}$. Then by following the analysis similar to Theorem 3.5 by utilizing the union bound and the Chernoff bound, we can prove that $\mathbb{P}[\Theta^{\mathrm{thres},1}_{i,T}\subseteq\hat{\Theta}^{\star}]\leq|\bar{\Theta}|e^{-2K\pi_{1}^{2}}$. Subsequently setting $K\geq\frac{\log(|\Theta^{\star}|/\eta_{1})}{2\pi_{1}^{2}}$ makes the error at most $\eta_{1}$.
Optimal Distributions. The proof is exactly the same as in Theorem 3.2.
Appendix B Online Learning
B.1. Log-Belief Ratio Notations
Let $\bm{\mu}_{i,t}(\hat{\theta})$ be the beliefs for the non-private system, and let $\bm{\nu}_{i,t}(\hat{\theta})$ be the private estimates for agent $i\in[n]$ and round $t\geq 0$. For the non-private algorithm (Algorithm 2) all pairs $\hat{\theta},\check{\theta}\in\Theta$ we let $\bm{\phi}_{i,t}(\hat{\theta},\check{\theta})=\log\left(\frac{\bm{\mu}_{i,t}(\hat{\theta})}{\bm{\mu}_{i,t}(\check{\theta})}\right)$ and $\bm{\lambda}_{i,t}(\hat{\theta},\check{\theta})=\log\left(\frac{\bm{\gamma}_{i,t}(\hat{\theta})}{\bm{\gamma}_{i,t}(\check{\theta})}\right)$ be the log belief ratios and the log-likelihood ratios respectively for agent $i\in[n]$ round $t\geq 0$. For the private algorithm (Algorithm 5) we let $\bm{\psi}_{i,t}(\hat{\theta},\check{\theta})=\log\left(\frac{\bm{\nu}_{i,t}(\hat{\theta})}{\bm{\nu}_{i,t}(\check{\theta})}\right),\bm{\zeta}_{i,t}(\hat{\theta},\check{\theta})=\log\left(\frac{\bm{\sigma}_{i,t}(\hat{\theta})}{\bm{\sigma}_{i,t}(\check{\theta})}\right),$ and $\bm{\kappa}_{i,t}(\hat{\theta},\check{\theta})=\bm{d}_{i,t}(\hat{\theta})-\bm{d}_{i,t}(\check{\theta})$ be the private log-belief ratio, log-likelihood ratio, and noise difference ratios respectively for agent $i\in[n]$, and round $t\geq 0$. We also let the vectorized versions $\overline{\bm{\phi}}_{t}(\hat{\theta},\check{\theta}),\overline{\bm{\psi}}_{t}(\hat{\theta},\check{\theta}),\overline{\bm{\lambda}}_{t}(\hat{\theta},\check{\theta}),\overline{\bm{\zeta}}_{t}(\hat{\theta},\check{\theta})$ respectively, where the vectorization is over the agents $i\in[n]$. Finally, for each agent $i\in[n]$ and pair of states $\hat{\theta},\check{\theta}\in\Theta$ we define the KL divergence between the states $\Lambda_{i}(\hat{\theta},\check{\theta})=\mathbb{E}_{\theta}\left[\log\left(\frac{\ell_{i}(\bm{s}_{i,0}|\hat{\theta})}{\ell_{i}(\bm{s}_{i,0}|\check{\theta})}\right)\right]=D_{KL}\left(\ell_{i}(\cdot|\hat{\theta})|\ell_{i}(\cdot|\check{\theta})\right)$.
B.2. Auxiliary Lemmas
Before proving the main result for online learning, we prove the following Lemma regarding the rate of convergence of the Césaro means.
Lemma B.1 ().
Let $\overline{\bm{X}}_{1},\dots,\overline{\bm{X}}_{t}$ be i.i.d. random variables, let $A$ be a doubly-stochastic matrix with largest magnitude of the non-trivial eigenvalue $\beta^{\star}_{n}=\max\{\lambda_{2}(A),|\lambda_{n}(A)|\}$ and let $\bm{x}_{\tau}={\frac{1}{n}\mathds{1}^{T}\overline{\bm{X}}_{\tau}}$, with $\max_{1\leq\tau\leq t}\sqrt{\mathbb{V}\left[\bm{x}_{\tau}\right]}\leq V$. Then
$\mathbb{E}\left[\left\|\frac{1}{t}\sum_{\tau=0}^{t-1}A^{t-\tau}\overline{\bm{X}}_{\tau}-\left(\frac{1}{t}\sum_{\tau=0}^{t-1}\bm{x}_{\tau}\right)\mathds{1}\right\|_{2}\right]\leq\frac{(n-1)V}{(1-(\beta^{\star}_{n})^{2})\sqrt{t}}.$
Proof.
We have that
$\mathbb{E}\left[\left\|A^{t-\tau}\overline{\bm{X}}_{\tau}-\bm{x}_{\tau}\mathds{1}\right\|_{2}\right]$ $=\mathbb{E}\left[\left\|\frac{1}{n}\mathds{1}\mathds{1}^{T}\overline{\bm{X}}_{\tau}+\sum_{i=2}^{n}\lambda_{i}(A)^{t-\tau}\overline{\bm{r}}_{i}\overline{\bm{l}}_{i}^{T}\overline{\bm{X}}_{\tau}-\bm{x}\mathds{1}\right\|_{2}\right]$
$=\mathbb{E}\left[\left\|\bm{x}_{\tau}\mathds{1}+\sum_{i=2}^{n}\lambda_{i}(A)^{t-\tau}\overline{\bm{r}}_{i}\overline{\bm{l}}_{i}^{T}\overline{\bm{X}}_{\tau}-\bm{x}_{\tau}\mathds{1}\right\|_{2}\right]$
$\leq\sum_{i=2}^{n}\left|\lambda_{i}(A)^{t-\tau}\right|\|\overline{\bm{r}}_{i}\|_{2}\|\overline{\bm{l}}_{i}\|_{2}\mathbb{E}\left[\|\overline{\bm{X}}_{\tau}\|_{2}\right]$
$\leq(n-1)(\beta^{\star}_{n})^{t-\tau}\mathbb{E}\left[\|\overline{\bm{X}}_{\tau}\|_{2}\right]$
$\leq(n-1)(\beta^{\star}_{n})^{t-\tau}\mathbb{E}\left[\|\overline{\bm{X}}_{0}\|_{2}\right]$
$\leq(n-1)(\beta^{\star}_{n})^{t-\tau}V.$
Therefore,
$\mathbb{E}\left[\left\|\frac{1}{t}\sum_{\tau=0}^{t-1}A^{t-\tau}\overline{\bm{X}}_{\tau}-\left(\frac{1}{t}\sum_{\tau=0}^{t-1}\bm{x}_{\tau}\right)\mathds{1}\right\|_{2}\right]\leq\frac{(n-1)V}{(1-(\beta^{\star}_{n})^{2})\sqrt{t}}.$
Since $\sum_{\tau=0}^{t}(\beta^{\star}_{n})^{t-\tau}\leq\sum_{\tau\geq 0}(\beta^{\star}_{n})^{\tau}=\frac{1}{1-(\beta^{\star}_{n})^{2}}$.
In the sequel, we show the main result for online learning
B.3. Proof of Theorem 3.7
Proof.
We let $\Xi_{n}=\mathbb{V}\left[\sum_{i=1}^{n}n_{i,\tau}\right]=\sum_{i=1}^{n}\xi_{i}$. We define the following sequence of “bad” events:
$B_{n,t}=\left\{\max_{i\in[n]}\left|\frac{1}{t}\sum_{\tau=0}^{t-1}n_{i,\tau}-\nu_{i}\right|\geq\sqrt{\frac{\Xi_{n}}{\eta t}}\right\}.$
By applying the union bound and Chebyshev’s inequality, we can show that $\mathbb{P}[B_{n,t}]\leq\eta$.
We condition on the “bad” event $B_{n,t}$ not happening. We can prove that, conditioned on $B_{n,t}^{c}$,
$V_{n,\Theta}^{\prime}=\max_{\bar{\theta}\in\bar{\Theta},\theta^{\star}\in\Theta^{\star}}\sqrt{\mathbb{V}\left[\frac{1}{nt}\sum_{i=1}^{n}\sum_{\tau=0}^{t-1}n_{i,\tau}\bm{\zeta}_{i,\tau}(\hat{\theta},\check{\theta})\bigg{|}B_{n,t}^{c}\right]}\leq\frac{\sqrt{2}}{n}(nQ_{n,\Theta}+V_{n,\Theta})\left(\max_{i\in[n]}\nu_{i}+\sqrt{\frac{\Xi_{n}}{\eta t}}\right).$
The dynamics of the log-belief ratio obey $\overline{\bm{\psi}}_{t}(\hat{\theta},\check{\theta})=A\overline{\bm{\psi}}_{t-1}(\hat{\theta},\check{\theta})+\overline{\bm{\zeta}}_{t}(\hat{\theta},\check{\theta})=\sum_{\tau=0}^{t-1}A^{t-\tau}\overline{\bm{\zeta}}_{\tau}(\hat{\theta},\check{\theta})$, where $\overline{\bm{\zeta}}_{\tau}(\hat{\theta},\check{\theta})=\overline{\bm{\lambda}}_{\tau}(\hat{\theta},\check{\theta})+\overline{\bm{\kappa}}_{\tau}(\hat{\theta},\check{\theta})$. We apply Lemma B.1, and get that
$\mathbb{E}\left[\left\|\frac{1}{t}\overline{\bm{\psi}}_{t}(\hat{\theta},\check{\theta})-\frac{1}{nt}\sum_{\tau=0}^{t-1}\sum_{i=1}^{n}n_{i,\tau}\bm{\zeta}_{i,\tau}(\hat{\theta},\check{\theta})\mathds{1}\right\|_{2}\bigg{|}B_{n,t}^{c}\right]$ $\leq\frac{(n-1)V_{n,\Theta}^{\prime}}{\sqrt{t}(1-(\beta^{\star}_{n})^{2})}.$
Moreover
$\mathbb{E}\left[\left\|\frac{1}{nt}\sum_{\tau=0}^{t-1}\sum_{i=1}^{n}n_{i,\tau}\bm{\zeta}_{i,\tau}(\hat{\theta},\check{\theta})-\frac{1}{n}\sum_{i=1}^{n}\nu_{i}\Lambda_{i}(\hat{\theta},\check{\theta})\right\|_{2}\bigg{|}B_{n,t}^{c}\right]\leq\frac{V_{n,\Theta}^{\prime}}{\sqrt{t}}.$
Therefore, by Markov’s inequality and the triangle inequality, we get that for every $z>0$
$\mathbb{P}\left[\left\|\frac{1}{t}\overline{\bm{\psi}}_{t}(\hat{\theta},\check{\theta})-\frac{1}{n}\sum_{i=1}^{n}\nu_{i}\Lambda_{i}(\hat{\theta},\check{\theta})\mathds{1}\right\|_{2}>z\bigg{|}B_{n,t}^{c}\right]\leq\frac{V_{n,\Theta}^{\prime}}{z(1-(\beta_{n}^{\star})^{2}))\sqrt{t}}.$
Therefore, we can show that by applying a union bound over $\Theta^{\star}$, we have that with probability $1-\eta$, for all $\bar{\theta}\in\bar{\Theta}$
$\log\bm{\nu}_{i,t}(\bar{\theta})\leq\bm{\psi}_{i,t}(\bar{\theta},\theta^{\star})\leq-\frac{t}{n}l_{n,\Theta}+\frac{|\Theta|V_{n,\Theta}^{\prime}\sqrt{t}}{2\eta(1-(\beta_{n}^{\star})^{2})}.$
To make the RHS the log-belief threshold at most some value $-\varrho$ for some $\varrho>0$, we require
$t\geq T=\mathcal{O}\left(\sqrt{\frac{n}{l_{n,\Theta}}\max\left\{\frac{|\Theta|(nQ_{n,\Theta}+V_{n,\Theta})(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}/\eta})}{\eta(1-(\beta_{n}^{\star})^{2})},\sqrt{\frac{l_{n,\Theta}\varrho}{n}}\right\}}\right).$
To determine $\varrho$, note that $\bm{\nu}_{i,t}(\bar{\theta})\leq e^{-\varrho}$ for all $\bar{\theta}\neq\theta^{\star}$, and, thus,
$\bm{\nu}_{i,t}(\theta^{\star})=1-\sum_{\bar{\theta}\neq\theta^{\star}}\bm{\nu}_{i,t}(\bar{\theta})\geq 1-(|\Theta|-1)e^{-\varrho}.$
To determine a valid value of $\varrho$ we require $1-(|\Theta|-1)e^{-\varrho}\geq e^{-\varrho}$ which yields $\varrho\geq\log(|\Theta|)$. Therefore, setting $\varrho=\log(|\Theta|)$ and subsequently
$T=\mathcal{O}\left(\sqrt{\frac{n}{l_{n,\Theta}}\max\left\{\frac{|\Theta|(nQ_{n,\Theta}+V_{n,\Theta})(\max_{i\in[n]}\nu_{i}+\sqrt{\Xi_{n}/\eta})}{\eta(1-(\beta_{n}^{\star})^{2})},\sqrt{\frac{l_{n,\Theta}\log(|\Theta|)}{n}}\right\}}\right),$
we get that
$\mathbb{P}\left[\theta^{\star}=\hat{\theta}^{\mathrm{OL}}_{i,T}\right]\geq\mathbb{P}[B_{n,T}^{c}]\mathbb{P}\left[\theta^{\star}=\hat{\theta}^{\mathrm{OL}}_{i,T}|B_{n,T}^{c}\right]\geq(1-\eta)^{2}\geq 1-2\eta.$
Privacy. Similarly to Theorem 3.2, the optimal distributions are the ones that minimize the variance subject to privacy constraints. Since there are $|\Theta|$ states, the budget ${\varepsilon}$ should be divided by $|\Theta|$, yielding $\mathcal{D}_{i,t}^{\star}({\varepsilon})=\mathrm{Lap}\left(\frac{\Delta_{i,t}|\Theta|}{{\varepsilon}}\right)$.
Remark. We note here that if we also want to protect the number of signals each agent is getting – thus extending the definition of Signal DP to the whole dataset – one should take the maximum over $\Delta_{i,t}$ and use it as sensitivity.